### v1

# Unsupervised Prior Learning: Discovering Categorical Pose Priors from Videos

Abstract
A prior represents a set of beliefs or assumptions about a system, aiding inference and decision-making. In this work, we introduce the challenge of unsupervised prior learning in pose estimation, where AI models learn pose priors of animate objects from videos in a self-supervised manner. These videos present objects performing various actions, providing crucial information about their keypoints and connectivity. While priors are effective in pose estimation, acquiring them can be difficult. We propose a novel method, named Pose Prior Learner (PPL), to learn general pose priors applicable to any object category. PPL uses a hierarchical memory to store compositional parts of prototypical poses, from which we distill a general pose prior. This prior enhances pose estimation accuracy through template transformation and image reconstruction. PPL learns meaningful pose priors without any additional human annotations or interventions, outperforming competitive baselines on both human and animal pose estimation datasets. Notably, our experimental results reveal the effectiveness of PPL using learnt priors for pose estimation on occluded images. Through iterative inference, PPL leverages priors to refine estimated poses, regressing them to any prototypical poses stored in memory. Our code, model, and data will be publicly available.
1 Introduction
Figure 1: Schematic illustration of the unsupervised prior learning challenge on videos in the task of pose estimation. Given a series of original video frames (framed in blue), the challenge is to learn the pose prior (green circle) in a self-supervised manner. The pose prior comprises keypoint and connectivity priors. To address this challenge, we propose a Pose Prior Learner (PPL). During PPL training, prototypical poses in the memory (blue circle) are aggregated from the individual poses estimated from all video frames (orange circle) and distilled into a general pose prior. The pose prior can then guide pose estimation from video frames through transformations. This cyclic process strengthens the learning of robust pose prior representations, resulting in more accurate pose estimations, which, in turn, helps capture more representative prototypical poses. During inference, the learned pose prior refines pose estimation on occluded images (bounded in red), regressing them to the prototypical poses stored in memory. Blue arrows illustrate signal flows during PPL training, while red arrows indicate signal flows during inference on occluded images.
Priors represent beliefs or assumptions about a system or the characteristics of a concept. They are widely used in statistical inference (Lindley, 1961), cognitive science (Schad et al., 2021), and machine learning (Diligenti et al., 2017; Gülçehre & Bengio, 2016). This pre-existing knowledge is essential for guiding the inference process, enabling AI models to make robust predictions in uncertain or ambiguous situations (Thiruvenkadam et al., 2008; Sung et al., 2015; Liang et al., 2024). The objective of our work is to enhance our understanding of priors in AI models and offer preliminary answers to the three key intelligence questions: (1) How do we acquire priors in the first place? (2) Can we learn them from input data in a self-supervised manner? (3) Can we evaluate and enhance the quality of the priors? To tackle these questions, we first introduce the challenge of unsupervised prior learning in the context of pose estimation from videos. See Figure 1 for the schematic illustration of the challenge. Pose estimation is a classical computer vision task that identifies the structure of objects in images and videos by detecting keypoints. A pose prior summarizes the common characteristics shared by a variety of poses. It encapsulates the expectation of the keypoint configurations and the connectivity between keypoints.
In parallel to our challenge of unsupervised prior learning from videos for pose estimation, unsupervised pose estimation leverages the abundant, unannotated visual information available in large image and video datasets to extract pose information (Hu & Ahuja, 2021; Sommer et al., 2024; Chen et al., 2019; He et al., 2022a; Schmidtke et al., 2021). The use of pose priors can provide valuable guidance in this process. We categorize the existing works into two groups: those that incorporate priors and those that operate without them.
Recent approaches (He et al., 2022a; Sun et al., 2022; 2023) attempt to predict keypoints from images, construct object structure representations using these keypoints, and learn effective structural information through image reconstruction. However, without pose priors, these methods can be easily disrupted by background information, causing the model to occasionally learn the background structure instead of the foreground object. While using video streams with static backgrounds helps alleviate this issue (Schmidtke et al., 2021; Yoo & Russakovsky, 2023), these methods remain vulnerable to producing inaccurate keypoints. This vulnerability arises from the lack of supplementary knowledge that could constrain both keypoint localization and the connectivity among keypoints.
In contrast, some methods (Schmidtke et al., 2021; Yoo & Russakovsky, 2023) utilize prior knowledge of a category’s pose to guide the pose estimation of individuals within that category. Conceptually, each category is expected to exhibit a generalized and distinctive pose prior that reflects characteristics such as shape, size, and structure. Individual poses should be seen as geometric transformations of this category-specific pose prior. As a result, employing a category-specific pose prior aids in guiding and regularizing the learning of poses. However, acquiring comprehensive general pose priors is highly challenging, as it necessitates extensive human annotations and domain knowledge.
Loosely inspired by how humans develop a universal prior representation of an object category by observing the movements of individual object instances in videos and subsequently using these priors to estimate individual poses, we propose a new method called the Pose Prior Learner (PPL). PPL is designed to effectively learn a meaningful pose prior for any object category. It utilizes a hierarchical memory to store a finite set of prototypical poses and extract a general pose prior from them. Initially, both the hierarchical memory and the universal prior representation are randomly initialized but learnable parameters. During training, effective pose learning is ensured through image reconstruction. As training progresses, the hierarchical memory retains and aggregates multiple accurate prototypical poses, thereby contributing to a more precise pose prior and enhancing the model’s ability to estimate poses.
Upon completing the training, we obtain a model that enables accurate pose estimation, a pose prior that encapsulates the general features of a category, and a hierarchical memory that stores diverse prototypical poses for that category. We evaluate the effectiveness of our PPL across several human and dog pose estimation benchmarks. We visualize their pose priors to further interpret what our approach has learned. Additionally, we introduce an iterative inference strategy to estimate the poses of objects in occluded scenes using the trained hierarchical memory and the pose prior. Our contributions are highlighted below:
1. We introduce the challenge of unsupervised prior learning in the context of pose estimation. We also establish evaluation metrics and benchmarks to assess the quality of the learned priors.
2. We propose a new method called Pose Prior Learner (PPL) for unsupervised pose estimation. PPL effectively mitigates background noise and does not require any domain knowledge from humans. The compositional hierarchical memory in PPL aggregates prototypical poses from training videos and subsequently utilizes these poses to distill pose priors. Three training techniques are introduced to ensure training convergence and stability of PPL.
3. PPL outperforms existing methods across several pose estimation benchmarks and offers explainable visualizations of pose priors. Notably, We found that predefined human priors are not always optimal. Our PPL even outperforms models using human-defined priors.
4. During inference, we utilize an iterative strategy in which PPL progressively leverages priors to refine estimated poses by regressing them to the nearest prototypical poses stored in memory. Experimental results demonstrate that our PPL accurately estimates poses, even in occluded scenes.
2 Related Works
Unsupervised Pose Estimation without Priors. Numerous methods without priors have been proposed to detect keypoints from images or videos, which are then used to reconstruct images for supervision (Li et al., 2021; Geng et al., 2021; Zhang et al., 2018; Sun et al., 2022; Thewlis et al., 2017; Jakab et al., 2020). For example, AutoLink (He et al., 2022a) extracts keypoints from the image and estimates the strength of the links between pairs of keypoints. It then combines these keypoints with the link heatmap to reconstruct the randomly masked image. Similarly, BKind (Sun et al., 2022; 2023) uses keypoints extracted from two video frames to reconstruct the pixel-level differences between these two frames. In these methods, keypoints are directly predicted from the image and supervised solely by image reconstruction, leading to potential detection of keypoints in background regions with complex textures. Moreover, the absence of constraints on keypoint configuration and connectivity also contributes to the unreliability of their methods. In contrast, our PPL utilizes the learned pose prior as a constraint to mitigate these issues.
Unsupervised Pose Estimation Incorporating Priors. Several methods utilize prior knowledge to guide the pose estimation (Chen & Dou, 2021; Shi et al., 2023; Zhang et al., 2022; Schmidtke et al., 2021; Yoo & Russakovsky, 2023). Among these methods, Shape Template Transforming (STT) (Schmidtke et al., 2021) applies affine transformations to a predefined pose prior, aligning it with the estimated pose from a video frame. By incorporating an additional frame from the same video to provide background information, an image reconstruction loss supervises the pose estimation process. The pose prior effectively guides pose estimation by constraining the shape of the human pose and the connectivity between body parts. However, pose priors are often difficult to obtain, requiring costly human annotations. Moreover, HPE (Yoo & Russakovsky, 2023) has shown that predefined pose priors are not always optimal, and tuning the shape of the prior can sometimes improve performance. Unlike these methods, our approach learns the prior directly from input videos, and models with our learned priors even outperform those using human-defined priors.
Compositional Memory Architectures. Compositional memory has been widely used in many computer vision tasks, such as question answering (Seong et al., 2021), object segmentation (Seong et al., 2021), and sence graph generation (Deng et al., 2022). In pose estimation, PCT (Geng et al., 2023) decomposes a human pose into discrete tokens, where each token connects several interdependent joints and characterizes a sub-structure of the entire human pose. This approach is highly effective for decomposing and reconstructing poses, providing robust pose representations while significantly reducing computation and storage costs. However, PCT encodes all tokens into the same embedding space, making it difficult to aggregate semantic tokens that represent different sub-structures. In contrast, our PPL employs a compositional hierarchical memory, which parses poses into memory banks. Each memory bank explicitly contains multiple tokens encoding the variations of each sub-structure of a pose. This facilitates the aggregation of poses into a universal pose prior.
3 Our Proposed Method – Pose Prior Learner (PPL)
We introduce our proposed method, Pose Prior Learner (PPL). Given videos with static backgrounds featuring object instances from a specific category, such as dogs or humans, PPL can accurately estimate the poses of the objects in that category while gradually learning a general pose prior through unsupervised learning. Note that our PPL requires no extra knowledge from human annotators. The architecture of PPL is presented in Figure 2.
Mathematically, We represent the topology for an object as a graph connecting keypoints with shared link weights, also known as connectivity. For any category, its general pose prior $V$ is defined as $V=(T,W)$, where $T$ represents its keypoint prior and $W$ denotes the connectivity prior. Specifically, $T$ consists of $N$ keypoints: $T=[P_{1},P_{2},...,P_{N}]$, where $P_{i}\in[-1,1]\times[-1,1]$ is the normalized pixel coordinates. $W$ is a $2D$ matrix of size $N\times N$, where each entry $w_{ij}$ in the matrix represents the connectivity probability between two keypoints $P_{i}$ and $P_{j}$. For instance, in the case of humans, the hand is connected to the torso via an arm; thus, the connectivity probability between these two parts should be higher than the connectivity probability between a hand and a foot. We randomly initialize the connectivity prior $W$ with positive values. The keypoint prior $T$ is decoded from a hierarchical memory $M$ storing compositional parts of prototypical poses.
Figure 2: Overview of our proposed Pose Prior Learner (PPL). We first distill the keypoint prior from the hierarchical memory $M$. Features of the frame $I$ and the embedding of the keypoint prior are concatenated to predict the affine transformation parameters. The keypoint prior is transformed and their pair-wise links are modulated with the connectivity prior $W$ to obtain the combined link heatmap $S$. The concatenation of the link heatmap $S$ and the reference frame $I_{ref}$ is decoded to produce the reconstructed frame $I_{recon}$. The $sg$ symbol represents the stopping gradient operation. The red arrows indicate the gradient flows during backpropagation based on image reconstruction. See Section 3.4 for training details.
During training, PPL inputs a pair of frames $I$ and $I_{ref}$ with the same sizes of $\mathsf{H}\times\mathsf{W}\times 3$ where $\mathsf{H}$ and $\mathsf{W}$ are the height and the width of each video frame. The reference frame $I_{ref}$ is a randomly selected frame from the same video, containing the same person as $I$ but in a different pose. The aim of PPL is to learn to correctly predict the keypoints $T^{\prime}$ and their connectivity on $I$. Ideally, if PPL makes perfect predictions of the pose on $I$, by combining it with the background information from $I_{\text{ref}}$, the reconstructed image $I_{\text{recon}}$ should match $I$ exactly. Next, we introduce how we estimate $T^{\prime}$ on $I$ using the keypoint prior $T$.
3.1 Transformation of the keypoint prior
Given the frame $I$, we use a feature extractor $\phi_{enc}$ to extract its embedding $h_{I}$: $h_{I}=\phi_{enc}(I)$. $\phi_{enc}$ is a 2D-Convolution Neural Network (2D-CNN) trained from scratch. The keypoint prior $T$ is converted into an embedding $h_{T}$ via a series of fully connected layers. Together with $h_{T}$ as inputs, PPL learns to predict the affine transformation parameters $\Theta_{i}\in[\Theta_{1},\Theta_{2},...,\Theta_{N}]$ for each keypoint $P_{i}$ in $T$ from $h_{I}$ via a two-layer fully connected network denoted as $FC(\cdot)$:
$[\Theta_{1},\Theta_{2},...,\Theta_{i},...,\Theta_{N}]=FC(h_{I},h_{T}),\text{ where }\Theta_{i}=\begin{bmatrix}a^{(i)}&b^{(i)}&t_{x}^{(i)}\\ c^{(i)}&d^{(i)}&t_{y}^{(i)}\\ 0&0&1\end{bmatrix},$ (1)
where $t_{x}^{(i)}$ and $t_{y}^{(i)}$ are the translations and $a^{(i)},b^{(i)},c^{(i)},d^{(i)}$ are the coefficients that define rotation, scaling, and shear. Each point $P_{i}$ in $T$ can then be transformed by $\Theta_{i}$, resulting in the keypoints $T^{\prime}$ for the frame $I$:
$T^{\prime}=[P^{\prime}_{1},P^{\prime}_{2},...,P^{\prime}_{N}],\text{ where }[P^{\prime}_{i},1]^{\top}=\Theta_{i}[P_{i},1]^{\top}.$ (2)
3.2 Connecting keypoints based on the connectivity prior
The connectivity of keypoints in objects is often fixed and rigid; for example, human arms maintain a relatively constant length, with a hand always connected to the torso via an arm and never connected to a foot. This rigidity in connectivity serves as a constraint, aiding in the regularization of pose estimation. In this section, we introduce the connectivity prior and explain how it can be used to regularize the connectivity strength between any pair of estimated keypoints in $T^{\prime}$ on $I$.
Similar to AutoLink (He et al., 2022a), PPL connects any two keypoints $P^{\prime}_{i}$ and $P^{\prime}_{j}$ in $T^{\prime}$ to obtain differentiable link heatmap $S_{i,j}\in\mathbb{R}^{\mathsf{H}\times\mathsf{W}}$. Intuitively, each 2D link heatmap represents a probability density map, where the pixel values along the link between two points are high, while other areas are assigned values close to zero. For any point $P^{\prime}_{i}$, its strongest connectivity to any of the other points in $T^{\prime}$ is activated on the combined link heatmap $S\in\mathbb{R}^{\mathsf{H}\times\mathsf{W}}$ via a max pooling operation over all the $N\times N$ link heatmaps:
$S=\max_{i,j}^{N\times N}(w_{i,j}S_{i,j}),$ (3)
where $w_{i,j}$ in the connectivity prior $W$ modulates the link heatmap $S_{i,j}$ based on whether the two keypoints $P^{\prime}_{i}$ and $P^{\prime}_{j}$ are physically connected. Ideally, if PPL correctly estimates the probability of physical links for an object category, $S_{i,j}$ will receive higher connectivity values, thereby activating the locations linking these two keypoints on the combined link map $S$.
Given the combined link map $S$ and the reference frame $I_{ref}$, PPL can reconstruct the image $I$. $I_{ref}$ shares the same background as $I$ but features a different foreground pose, thereby offering background information for reconstruction. Meanwhile, $S$ provides the connectivity details among all the estimated keypoints. This encourages PPL to learn accurate pose estimations while minimizing the influence of background noise. Therefore, we concatenate $S$ and $I_{ref}$ and feed them into a 2D-CNN to perform the image reconstruction $I_{recon}$, where $I_{recon}=\phi_{dec}(I_{ref},S)$.
3.3 Reconstruction of Keypoint Configuration with Memory
$M$ is a hierarchical memory storing compositional representations of prototypical poses. It contains $m$ memory banks $\{b_{1},b_{2},...,b_{m}\}$ of $k$ learnable tokens with $d$ randomly initialized parameters per token. $M$ is important to learn robust prior representations $V$ due to the following reasons: (1) By storing multiple prototypical poses, memory helps in aggregating these poses, which aids in creating a more robust and comprehensive prior that captures variations within an object category. (2) $M$ can assist in organizing information hierarchically, making it easier for PPL to retrieve relevant prototypical poses when making predictions in uncertain or ambugious scenarios, such as occlusion. and (3) By leveraging $M$, PPL can refine its predictions iteratively, using stored poses to adjust its outputs based on previously learned pose representations. Next, we introduce how $M$ is structured.
Given the estimated $N$ keypoints in $T^{\prime}$, we first encode it into $m$ tokens of dimension $d$ with several MLP-Mixer blocks (Tolstikhin et al., 2021), denoted as $MIX_{enc}$. We define $G$ as the collection of the $m$ tokens, where each token $g_{i}$ represents the embedding of a subset of the keypoints. Each token $g_{i}$ shares the same embedding space with the memory bank $b_{i}$ in memory $M$:
$G=[g_{1},g_{2},...,g_{m}]=MIX_{enc}(T^{\prime}),$ (4)
If all the tokens in all the memory banks of $M$ learn to capture unique parts of prototypical poses, each $g_{i}$ should be able to retrieve the most similar token in the memory bank $b_{i}$ of $M$ and reconstructs $G$ itself. Here, we use $L2$ distance to measure the similarity between $g_{i}$ and $k$ token in $b_{i}$ of $M$. The token in $b_{i}$ that is most similar to $g_{i}$ is denoted as $g^{\prime}_{i}$. Thus, for each $g_{i}$, we can always find the corresponding $g^{\prime}_{i}$ from $b_{i}$, resulting in a collection of $G^{\prime}=[g^{\prime}_{1},g^{\prime}_{2},...,g^{\prime}_{m}]$. Another series of MLP-Mixer blocks (denoted as $MIX_{dec}$) is then used to decode $G^{\prime}$ back to $N$ keypoints $T^{\prime}_{recon}$, where $T^{\prime}_{recon}=MIX_{dec}(G^{\prime})$. See Figure LABEL:fig:mem_stru for the schematic.
Unlike PCT (Geng et al., 2023) where all tokens of a memory are within the same embedding space, We organize $M$ into $m$ memory banks, each containing $k$ tokens that represent an independent embedding space. This structure enables us to efficiently distill compositional parts from each memory bank, allowing us to form a general representation of a pose, which is distilled into the keypoint prior $T$. Specifically, we define $MP(\cdot)$ as the mean pooling operation. PPL pools one token vector among the $k$ vectors of each memory bank. These pooled $m$ vectors are further decoded by $MIX_{dec}$ into $N$ points of our distilled keypoint prior $T$. See Figure LABEL:fig:mem_cons for the schematic of distilling $T$: $T=MIX_{dec}([MP(b_{1}),MP(b_{2}),...,MP(b_{m})])$.
3.4 Training and Inference
Our PPL is trained to jointly minimize all the four losses: the image reconstruction loss $L_{ir}$, the boundary loss $L_{b}$, the link regularization loss $L_{l}$, and the keypoint configuration reconstruction loss $L_{kr}$. We elaborate on these four losses below.
Image Reconstruction Loss. If PPL correctly estimates the pose on the original image $I$, the reconstructed image $I_{recon}$, based on the estimated pose, should be identical to $I$. Therefore, ensuring the quality of $I_{recon}$ encourages PPL to improve its pose estimation accuracy. To achieve this, we apply a perceptual loss on the embeddings of $I$ and $I_{recon}$, extracted using a frozen feature extractor $\psi(\cdot)$ from the VGG19 network pre-trained on ImageNet (Russakovsky et al., 2015). The perceptual loss is defined as: $L_{ir}=\|\psi(I_{recon})-\psi(I)\|_{1}$.
Boundary Loss. To ensure that the network does not transform the points in the keypoint prior outside the boundaries of the image, we limit the x and y coordinates of the transformed keypoints to be within the image:
$L_{b}=\sum_{*\in{x,y}}\begin{cases}|P^{\prime}_{i,*}|&\text{if }|P^{\prime}_{i,*}|>1,\\ 0&\text{otherwise}.\end{cases}\text{ and }L_{l}=\sum_{i,j}w_{i,j}\|l(P_{i},P_{j})-l(P^{\prime}_{i},P^{\prime}_{j})\|_{1},$ (5)
where $P^{\prime}_{i,x}$ and $P^{\prime}_{i,y}$ are the normalized $x$ and $y$ coordinate of the keypoint $P^{\prime}_{i}$ respectively.
Link Regularization Loss. A person’s arm always maintains a fixed length regardless of the poses. Thus, we propose the constraint that links should be assigned a high weight if they do not vary significantly in length before and after the affine transformation. The loss $L_{l}$ encourages the preservation of link lengths during pose estimation. It is defined as in the equation above, where $l(\cdot)$ is the $L2$ distance between two keypoints before and after the affine transformation.
Reconstruction Loss on Keypoint Configurations. In Section 3.3, given a collection of token representations in $G$, PPL retrieves the most similar tokens from each memory bank of the hierarchical memory $M$ and generates $G^{\prime}$ in a non-differentiable manner. To ensure that $M$ learns to store meaningful token embeddings that represent compositional parts of poses, the retrieved token representations $G^{\prime}$ from $M$ should closely match $G$. Moreover, if these compositional parts are structured correctly, the tokens should be able to decode into meaningful keypoint configurations $T^{\prime}_{recon}$ that are close to the original keypoint configurations $T^{\prime}$. Therefore, we introduce the keypoint configuration reconstruction loss $L{kr}$, defined as: $L_{kr}=\|T^{\prime}_{recon}-T^{\prime}\|_{2}+\|G-G^{\prime}\|_{2}$.
Training Techniques. To ensure convergence and stability during the training of PPL, we introduce three gradient dettachment techniques: (1) To address the broken gradient issue during the quantization step from $G$ to $G^{\prime}$, we adopt the approach from VQ-VAE (Van Den Oord et al., 2017). Specifically, our PPL copies the gradients of $G^{\prime}$ to $G$ for backward propagation, allowing the gradients to flow through the quantization step. (2) The hierarchical memory $M$ is updated using an exponential moving average to smooth the gradient updates, particularly during the early stages of training when $G$ can be quite noisy. This approach helps stabilize the learning process and ensures that $M$ retains more reliable information over time. (3) For $M$ to learn effective representations of $G$, it requires an accurate estimation of $T^{\prime}$, which depends on a good prior $V$ distilled from $M$. This creates a chicken-and-egg problem that complicates training. To address this, we introduce two gradient detachments to separate the training processes. First, we detach the gradients from $T$ and train the keypoint transformation and image reconstruction pathway, as shown by the red arrows in Figure 2. Second, we detach the gradients from $T^{\prime}$ to train the memory encoder and decoder, $MIX_{\text{enc}}$ and $MIX_{\text{dec}}$, as indicated by the green arrows in Figure LABEL:fig:mem_stru.
Iterative Inference. We propose an iterative inference strategy (Figure 4). $4$ iterations are used for every experiment. In every iteration, we take the reconstructed frame $I_{recon}$ from the last iteration (the original frame $I$ for iteration 00) as the input. We infer its keypoints $T^{\prime}$ as the output keypoints of the current iteration. The hierarchical memory $M$ is used to reconstruct $T^{\prime}$ and the reconstructed keypoints $T^{\prime}_{recon}$ are used to obtain the reconstructed frame $I_{recon}$. $I_{recon}$ is then used as the input for the next iteration. We keep the original occluded frame $I$ as the reference frame for all the iterations. See Appendix A1 for more implementation details.
Figure 4: Overview of the iterative inference strategy in our PPL. During inference, we iteratively use the reconstructed frame $I_{recon}$ as input to estimate the pose $T^{\prime}$. The hierarchical memory $M$ refines the estimated pose $T^{\prime}$ and outputs $T^{\prime}_{recon}$. With the original test image $I$ as the reference frame $I_{ref}$, PPL reconstructs the image $I_{recon}$. It is then used as the input frame in the next iteration. See Section 3.4 for details.
4 Experiments
We use three video datasets: Human3.6m (Ionescu et al., 2013), Taichi (Siarohin et al., 2019), and YouTube dog videos with green backgrounds for our experiments. See Appendix A2 for details. On Human3.6m, we report the results in the mean $L2$ error between the predicted keypoints and the ground truth, normalized by the image size. For Taichi, as in (He et al., 2022a; Siarohin et al., 2021; He et al., 2021; 2022b; Zhang et al., 2018) we use the summed $L2$ error computed at a resolution of $256\times 256$. To ensure a more accurate evaluation, unlike previous methods (He et al., 2022a; Schmidtke et al., 2021), we employ a minimum flow algorithm (Waissi, 1994) to align the predicted keypoints with the ground truth.
Table 1: Keypoint detection on Human3.6m. We report the mean $L2$ error normalized by the image resolutions (Res.) of both $256\times 256$ and $128\times 128$. * is the $L2$ error calculated by our minimum flow algorithm. The best is in bold.
Method Res. Norm. $L2$ Error
(Jakab et al., 2020) 256 2.73
(Thewlis et al., 2017) 256 7.51
(Lorenz et al., 2019) 256 2.79
(Zhang et al., 2018) 256 4.91
(Schmidtke et al., 2021) 256 3.31
AutoLink (He et al., 2022a) 128 2.76
AutoLink* (He et al., 2022a) 128 2.61
PPL (ours) 128 1.92
PPL (ours) 256 2.56
Table 2: Keypoint detection on Taichi. For consistency with the baseline methods, we report the summed $L2$ error at a resolution of $256\times 256$. * is the $L2$ error calculated by our minimum flow algorithm. The best is in bold.
Method Summed $L2$ Error
(Siarohin et al., 2021) 389.78
(He et al., 2021) 437.69
(Zhang et al., 2018) 343.67
(He et al., 2022b) 417.17
AutoLink (He et al., 2022a) 316.10
AutoLink* (He et al., 2022a) 302.41
PPL (ours) 293.35
4.1 Unsupervised Human Pose Estimation
We compare the keypoint detection results of PPL with other unsupervised pose estimation methods and present the results in Table 1 (Human3.6m) and Table 2 (Taichi). On both datasets, PPL significantly outperforms all baselines across all image resolutions. Among the baselines, AutoLink (He et al., 2022a) also incorporates learnable connectivity priors, similar to our PPL. However, its performance is inferior due to the absence of hierarchical memory. Additionally, we note that STT (Schmidtke et al., 2021), a baseline with human-defined priors, still underperforms compared to our PPL. Consistent with HPE (Yoo & Russakovsky, 2023), this suggests that pre-defined priors are not always optimal, and our PPL is able to learn more representative priors that outperform those manually defined.
Visualization of Pose Estimation. We provide the visualization results of pose estimation on Human3.6m (Figure LABEL:fig:human-pose-vis) and YouTube dog videos (Figure LABEL:fig:dog-pose-vis). The visualization results demonstrate that PPL can learn priors and estimate poses of various object categories, without any external knowledge. For example, in Row 3, Column 2 of Figure LABEL:fig:human-pose-vis, PPL correctly estimates the bowing pose of a person. In Row 3, Column 5 of Figure LABEL:fig:dog-pose-vis, PPL correctly estimates the pose of a dog lowering down its head. Moreover, we found the quality of estimated poses of dogs is inferior to that of humans. This is primarily due to the significant morphological differences among various dog breeds. Additionally, dogs often perform actions such as turning around, which can lead to significant changes in pose that are difficult to accurately capture by our priors in 2D space.
Visualization of the Pose Prior Changing with the Training Epochs. We visualize the progressively learnt pose priors by our PPL as a function of training epochs. Figure LABEL:fig:prior-vis illustrates that the keypoint prior converges to a human shape by the early stage of training (epoch 5). Notably, the learnable keypoints align with the human joints defined in the literature, and the connectivity among keypoints corresponds to the physical connections between body parts. As training continues, the connectivity prior gradually learns the skeletal structure of the human body, with irrelevant links between keypoints diminishing over time (as seen when comparing epochs 15 and 20).
4.2 Ablation Studies
Table 3: Keypoint Detection Results of our PPL Variants on the Human3.6m dataset. All results in mean $L2$ errors are normalized by the image resolution of $256\times 256$. Both keypoint prior (Row 1-2) and Connectivity prior (Row 3-4) can be either pre-defined (Pre.) or randomly initialized (Rand.). During training, the parameters in both the priors can be either frozen (✗) or learnable (✓). The last column (From Mem) shows the result of our default PPL method. Its keypoint prior is initialized from memory (From Mem). Its connectivity prior is randomly initialized (Rand.) and learnable (✓) during training. Best is in bold.
1 2 3 4 5 6 7 8 9 10 11
Keypoint prior Initialization Pre Pre Pre Pre Pre Pre Rand Rand Rand Rand From mem
Trainable
Connectivity prior Initialization Pre Pre Pre Pre Rand Rand Pre Pre Rand Rand Rand
Trainable
Normalized $L2$ Error 2.51 2.66 2.58 2.70 2.54 2.61 2.68 2.72 2.75 2.83 2.56
Ablation on Prior Variants. Here, we investigate how different initializations of connectivity and keypoint priors affect pose estimation and assess whether further refining these priors enhances performance. From Table 3, we obtain several key insights: (1) Models with frozen, human-defined priors (Column 4) perform worse than our PPL, indicating that PPL learns more representative priors than those predefined by humans. (2) Refining pre-defined keypoint and connectivity priors (Column 1) outperforms our default PPL, suggesting that PPL can enhance models with human-defined priors through refinement. (3) Interestingly, randomly initializing either keypoint or connectivity priors, followed by refinement during training (Columns 5-9), yields comparable performance to models with human-defined priors. This suggests that human-defined priors may not be necessary for effective pose estimation. (4) Surprisingly, freezing randomly initialized keypoint priors also results in reasonable pose estimation accuracy, though it is still lower than PPL’s default performance (Columns 7 and 9). (5) In contrast to (4), freezing random connectivity priors prevents the model from converging, implying that connectivity priors play a more critical role in guiding pose estimations than keypoint priors.
Ablation on Number & Dimension of Tokens in Each Memory Bank. In our hierarchical memory, we used 34 memory banks. Here, we analyze the impact of the number of tokens per memory bank and the dimension of each token on PPL’s pose estimation performance. From Figure A1 in Appendix A3, we observed that PPL remains robust across different token counts and dimensions, although performance slightly improves with more tokens of higher dimensions. As a result, we fixed 16 tokens per memory bank, each with a dimension of 512, for all experiments.
Ablation on Number of Keypoints. We varied the number of keypoints in the pose priors from 4 to 32. The results in Figure A1 in Appendix A3 show that pose estimation accuracy improves as the number of keypoints in the prior increases. However, using 32 keypoints offers limited improvement compared to PPL with 16 keypoints for human pose estimation.
4.3 Pose Estimation in Occlusion Scene
To verify the robustness of PPL in occluded scenes, we divide the image into $32\times 32$ patches and apply two masking techniques: RandomMasking and CenterMasking. In RandomMasking, we randomly mask a certain proportion of image patches, with the proportion ranging from 0.1 to 0.4. In CenterMasking, we mask only the center region of the image, gradually increasing the masking size from $4\times 4$ to $12\times 12$ patches.
We explore the effect of occluded areas on PPL. From Figure 9 and Figure 9, we observe that at iteration 0, as the occluded areas increase, overall performance declines with larger occlusions. However, with our iterative inference strategy, PPL effectively infers the missing parts of the poses by utilizing prototypical poses stored in hierarchical memory and the learned priors. Notably, it restores partially occluded poses to reasonably complete full-body poses, leading to a lower $L2$ error, comparable to those without occlusion. This effect is more pronounced when the occlusion area is small or medium.
Visualization of Pose Estimation with Occlusion. We present the estimated poses by our PPL for occluded images as a function of the number of inference iterations in Figure LABEL:fig:occlusion-vis. Across both RandomMasking (top 2 rows) and CenterMasking (bottom 2 rows), with our iterative inference strategy, PPL successfully reconstructs the occluded part of the image after three iterations and meanwhile, predicts reasonable complete full-body poses.
Figure 8: PPL Results of Keypoint Detection as a Function of Number of Inference Iterations on Images with RandomMasking from Human3.6m. The “mask ratio” in the legend specifies the masked proportion on the $32\times 32=1024$ image patches.
Figure 9: PPL Results of Keypoint Detection as a Function of Number of Inference Iterations on Images with CenterMasking from Human3.6m. The “mask size” in the legend refers to the width and height of the masked region, on $1024$ image patches.
5 Discussion
We introduce the challenge of unsupervised prior learning and highlight its significance in pose estimation. To address this, we propose a novel method called Pose Prior Learner (PPL). PPL utilizes a hierarchical memory to store compositional parts of learnable prototypical poses, which are distilled into a general pose prior for any object category. Our experimental results show that PPL requires no additional knowledge and outperforms recent competitive baselines in video pose estimation. Notably, the learned prior proves to be even more effective in pose estimation than methods that rely on human-defined priors. With hierarchical memory and learned priors, PPL can perform iterative inferences and robustly estimate poses in occluded scenes.
Despite outstanding performance in unsupervised pose estimation, PPL has several limitations. For instance, it requires video data captured in static backgrounds. Additionally, PPL learns 2D priors, which makes it difficult to capture real-world 3D postures. Thus, PPL struggles in scenarios where foreground objects involve rotations or significant shape changes. Extending PPL to incorporate 3D priors will be a key focus of our future research.
Appendix A1 Implementation Details of our Pose Prior Learner
We use the Adam optimizer with a learning rate of $10^{-3}$ and a batch size of 64, training for 50 epochs. Unless specified, all images are resized to $256\times 256$. The learning rate for link weights is scaled by 512 to address the small gradients of SoftPlus near zero. We conduct experiments using the link thickness $\sigma^{2}=5\times 10^{-4}$ across all benchmark datasets, where we adopt the same definition of $\sigma$ used in (He et al., 2022a). For the hierarchical memory, we use $34$ memory banks, each of which contains $16$ tokens of dimension $512$, for all experiments.
Appendix A2 Datasets
Human3.6m (Ionescu et al., 2013) is a standard benchmark dataset for human pose estimation, consisting of 3.6 million video frames. These frames include both 3D and 2D keypoints and were captured in a controlled studio environment with a static background, featuring various actors. We adhere to the approach outlined in (Zhang et al., 2018; He et al., 2022a), focusing on six activities: direction, discussion, posing, waiting, greeting, and walking. For training, we use subjects 1, 5, 6, 7, 8, and 9, and for testing, we use subject 11.
Taichi (Siarohin et al., 2019) consists of 3,049 training videos and 285 test videos featuring individuals performing Tai-Chi, with diverse foreground and background appearances. Following the approach in (Siarohin et al., 2021), we use 5,000 frames for training and 300 frames for testing.
YouTube dog videos are videos with green backgrounds collected from YouTube to further qualitatively demonstrate the performance of PPL in non-human species. We used only 20 videos from YouTube and decomposed them into a total of 2000 video frames for training. We use this customized dataset to demonstrate the applicability of PPL in learning pose priors for non-human categories. All images are trained and tested at a resolution of $256\times 256$. We use $10$ keypoints for this category and provide the visualization of the estimated poses on the test videos.
Appendix A3 Ablation on Memory Bank Tokens and Number of Keypoints
Figure A1: Ablation of our PPL method on Memory Bank Tokens and Number of Keypoints in Human3.6m. The upper horizontal axis is the number of keypoints (ranging from 4 to 32) and the lower horizontal axis is the dimension of memory bank tokens (ranging from 64 to 512). The dashed purple line is for ablations on number of keypoints and the solid lines are for ablations on memory bank tokens.

### v2

@@
-Unsupervised Prior Learning: Discovering Categorical Pose Priors from Videos
+Pose Prior Learner: Unsupervised Categorical Prior Learning for Pose Estimation
 Abstract
 A prior represents a set of beliefs or assumptions about a system, aiding inference and decision-making.
-In this work, we introduce the challenge of unsupervised prior learning in pose estimation, where AI models learn pose priors of animate objects from videos in a self-supervised manner.
-These videos present objects performing various actions, providing crucial information about their keypoints and connectivity.
-While priors are effective in pose estimation, acquiring them can be difficult.
-We propose a novel method, named Pose Prior Learner (PPL), to learn general pose priors applicable to any object category.
+In this paper, we introduce the challenge of unsupervised categorical prior learning in pose estimation, where AI models learn a general pose prior for an object category from images in a self-supervised manner.
+Although priors are effective in estimating pose, acquiring them can be difficult.
+We propose a novel method, named Pose Prior Learner (PPL), to learn a general pose prior for any object category.
 PPL uses a hierarchical memory to store compositional parts of prototypical poses, from which we distill a general pose prior.
-This prior enhances pose estimation accuracy through template transformation and image reconstruction.
+This prior improves pose estimation accuracy through template transformation and image reconstruction.
 PPL learns meaningful pose priors without any additional human annotations or interventions, outperforming competitive baselines on both human and animal pose estimation datasets.
-Notably, our experimental results reveal the effectiveness of PPL using learnt priors for pose estimation on occluded images.
-Through iterative inference, PPL leverages priors to refine estimated poses, regressing them to any prototypical poses stored in memory.
+Notably, our experimental results reveal the effectiveness of PPL using learned prototypical poses for pose estimation on occluded images.
+Through iterative inference, PPL leverages the pose prior to refine estimated poses, regressing them to any prototypical poses stored in memory.
 Our code, model, and data will be publicly available.
 1 Introduction
-Figure 1: Schematic illustration of the unsupervised prior learning challenge on videos in the task of pose estimation.
-Given a series of original video frames (framed in blue), the challenge is to learn the pose prior (green circle) in a self-supervised manner.
+Figure 1: Schematic illustration of the unsupervised categorical prior learning challenge in the task of pose estimation.
+Given a series of images (framed in blue), the challenge is to learn the pose prior (green circle) in a self-supervised manner.
 The pose prior comprises keypoint and connectivity priors.
 To address this challenge, we propose a Pose Prior Learner (PPL).
-During PPL training, prototypical poses in the memory (blue circle) are aggregated from the individual poses estimated from all video frames (orange circle) and distilled into a general pose prior.
-The pose prior can then guide pose estimation from video frames through transformations.
+During PPL training, prototypical poses in the memory (blue circle) are aggregated from the individual poses estimated from all images (orange circle) and distilled into a general pose prior.
+The pose prior can then guide pose estimation through transformations.
 This cyclic process strengthens the learning of robust pose prior representations, resulting in more accurate pose estimations, which, in turn, helps capture more representative prototypical poses.
@@
 (2) Can we learn them from input data in a self-supervised manner?
-(3) Can we evaluate and enhance the quality of the priors?
-To tackle these questions, we first introduce the challenge of unsupervised prior learning in the context of pose estimation from videos.
+(3) Can we enhance the quality of the priors?
+To tackle these questions, we first introduce the challenge of unsupervised categorical prior learning in the context of pose estimation from images.
 See Figure 1 for the schematic illustration of the challenge.
-Pose estimation is a classical computer vision task that identifies the structure of objects in images and videos by detecting keypoints.
+Categorical pose estimation is a classical computer vision task that identifies the structure of objects belonging to the same category by detecting their keypoints.
 A pose prior summarizes the common characteristics shared by a variety of poses.
 It encapsulates the expectation of the keypoint configurations and the connectivity between keypoints.
-In parallel to our challenge of unsupervised prior learning from videos for pose estimation, unsupervised pose estimation leverages the abundant, unannotated visual information available in large image and video datasets to extract pose information (Hu & Ahuja, 2021; Sommer et al., 2024; Chen et al., 2019; He et al., 2022a; Schmidtke et al., 2021).
+In parallel to our challenge of unsupervised categorical prior learning from images for pose estimation, unsupervised pose estimation leverages the abundant, unannotated visual information available in large image datasets to extract pose information (Hu & Ahuja, 2021; Sommer et al., 2024; Chen et al., 2019; He et al., 2022a; Schmidtke et al., 2021).
 The use of pose priors can provide valuable guidance in this process.
-We categorize the existing works into two groups: those that incorporate priors and those that operate without them.
-Recent approaches (He et al., 2022a; Sun et al., 2022; 2023) attempt to predict keypoints from images, construct object structure representations using these keypoints, and learn effective structural information through image reconstruction.
-However, without pose priors, these methods can be easily disrupted by background information, causing the model to occasionally learn the background structure instead of the foreground object.
-While using video streams with static backgrounds helps alleviate this issue (Schmidtke et al., 2021; Yoo & Russakovsky, 2023), these methods remain vulnerable to producing inaccurate keypoints.
-This vulnerability arises from the lack of supplementary knowledge that could constrain both keypoint localization and the connectivity among keypoints.
-In contrast, some methods (Schmidtke et al., 2021; Yoo & Russakovsky, 2023) utilize prior knowledge of a category’s pose to guide the pose estimation of individuals within that category.
+We categorize the existing works in unsupervised pose estimation into two groups: those that incorporate hand-made priors and those that operate without any priors.
+Recent approaches (He et al., 2022a; Sun et al., 2022, 2023) attempt to predict keypoints from images, construct object structure representations using these keypoints, and learn effective structural information through image reconstruction.
+However, without pose priors, these methods can be disrupted by background information or may predict infeasible topological configurations of an object during occlusion.
+The risk of generating inaccurate keypoints stems from the absence of supplementary information that could help refine both keypoint localization and the connections between them.
+The other group of methods (Schmidtke et al., 2021; Yoo & Russakovsky, 2023) utilize prior knowledge of a category’s general pose to guide the pose estimation of individuals within that category.
 Conceptually, each category is expected to exhibit a generalized and distinctive pose prior that reflects characteristics such as shape, size, and structure.
@@
 As a result, employing a category-specific pose prior aids in guiding and regularizing the learning of poses.
-However, acquiring comprehensive general pose priors is highly challenging, as it necessitates extensive human annotations and domain knowledge.
-Loosely inspired by how humans develop a universal prior representation of an object category by observing the movements of individual object instances in videos and subsequently using these priors to estimate individual poses, we propose a new method called the Pose Prior Learner (PPL).
-PPL is designed to effectively learn a meaningful pose prior for any object category.
+However, obtaining comprehensive general pose priors is highly challenging, as it requires extensive human annotations, particularly for novel categories.
+Moreover, human annotations may introduce implicit biases, hindering models from learning more meaningful priors.
+Loosely inspired by how humans develop a general prior representation of an object category by observing individual object instances in images and subsequently using them to infer upcoming individual poses, we propose a new method called the Pose Prior Learner (PPL).
+PPL is designed to effectively learn a meaningful pose prior for a certain object category.
 It utilizes a hierarchical memory to store a finite set of prototypical poses and extract a general pose prior from them.
-Initially, both the hierarchical memory and the universal prior representation are randomly initialized but learnable parameters.
-During training, effective pose learning is ensured through image reconstruction.
+Initially, both the hierarchical memory and the prior are randomly initialized but learnable parameters.
+During training, effective pose learning is supervised through image reconstruction.
 As training progresses, the hierarchical memory retains and aggregates multiple accurate prototypical poses, thereby contributing to a more precise pose prior and enhancing the model’s ability to estimate poses.
-Upon completing the training, we obtain a model that enables accurate pose estimation, a pose prior that encapsulates the general features of a category, and a hierarchical memory that stores diverse prototypical poses for that category.
-We evaluate the effectiveness of our PPL across several human and dog pose estimation benchmarks.
+Upon completing the training, we obtain a model that enables accurate pose estimation, a categorical pose prior that encapsulates the general features of a category, and a hierarchical memory that stores diverse prototypical poses for that category.
+We evaluate the effectiveness of our PPL across several human and animal pose estimation benchmarks.
 We visualize their pose priors to further interpret what our approach has learned.
@@
 Our contributions are highlighted below:
-1. We introduce the challenge of unsupervised prior learning in the context of pose estimation. We also establish evaluation metrics and benchmarks to assess the quality of the learned priors.
+1. We introduce the challenge of unsupervised categorical prior learning in the context of pose estimation.
 2.
 We propose a new method called Pose Prior Learner (PPL) for unsupervised pose estimation.
-PPL effectively mitigates background noise and does not require any domain knowledge from humans.
-The compositional hierarchical memory in PPL aggregates prototypical poses from training videos and subsequently utilizes these poses to distill pose priors.
-Three training techniques are introduced to ensure training convergence and stability of PPL.
-3.
 PPL outperforms existing methods across several pose estimation benchmarks and offers explainable visualizations of pose priors.
@@
 Our PPL even outperforms models using human-defined priors.
-4.
+3.
 During inference, we utilize an iterative strategy in which PPL progressively leverages priors to refine estimated poses by regressing them to the nearest prototypical poses stored in memory.
@@
 Unsupervised Pose Estimation without Priors.
-Numerous methods without priors have been proposed to detect keypoints from images or videos, which are then used to reconstruct images for supervision (Li et al., 2021; Geng et al., 2021; Zhang et al., 2018; Sun et al., 2022; Thewlis et al., 2017; Jakab et al., 2020).
+Numerous unsupervised learning methods without priors have been proposed to detect keypoints from images, which are then used to reconstruct images for supervision (Li et al., 2021; Geng et al., 2021; Zhang et al., 2018; Sun et al., 2022; Thewlis et al., 2017; Jakab et al., 2020).
 For example, AutoLink (He et al., 2022a) extracts keypoints from the image and estimates the strength of the links between pairs of keypoints.
 It then combines these keypoints with the link heatmap to reconstruct the randomly masked image.
-Similarly, BKind (Sun et al., 2022; 2023) uses keypoints extracted from two video frames to reconstruct the pixel-level differences between these two frames.
 In these methods, keypoints are directly predicted from the image and supervised solely by image reconstruction, leading to potential detection of keypoints in background regions with complex textures.
-Moreover, the absence of constraints on keypoint configuration and connectivity also contributes to the unreliability of their methods.
+To alleviate this problem, BKind (Sun et al., 2022, 2023) uses keypoints extracted from two video frames to reconstruct the pixel-level differences between these two frames.
+However, the lack of constraints on keypoint configuration and connectivity still undermines the reliability of their approach.
 In contrast, our PPL utilizes the learned pose prior as a constraint to mitigate these issues.
-Unsupervised Pose Estimation Incorporating Priors.
-Several methods utilize prior knowledge to guide the pose estimation (Chen & Dou, 2021; Shi et al., 2023; Zhang et al., 2022; Schmidtke et al., 2021; Yoo & Russakovsky, 2023).
+Unsupervised Pose Estimation Incorporating Human-defined Priors.
+Several methods utilize prior knowledge from human annotators to guide the pose estimation (Chen & Dou, 2021; Shi et al., 2023; Zhang et al., 2022; Schmidtke et al., 2021; Yoo & Russakovsky, 2023).
 Among these methods, Shape Template Transforming (STT) (Schmidtke et al., 2021) applies affine transformations to a predefined pose prior, aligning it with the estimated pose from a video frame.
@@
 Moreover, HPE (Yoo & Russakovsky, 2023) has shown that predefined pose priors are not always optimal, and tuning the shape of the prior can sometimes improve performance.
-Unlike these methods, our approach learns the prior directly from input videos, and models with our learned priors even outperform those using human-defined priors.
+Unlike these methods, our approach learns the prior directly from input images without any manual annotations, and models with our learned priors even outperform those using human-defined priors.
 Compositional Memory Architectures.
@@
 In contrast, our PPL employs a compositional hierarchical memory, which parses poses into memory banks.
-Each memory bank explicitly contains multiple tokens encoding the variations of each sub-structure of a pose.
-This facilitates the aggregation of poses into a universal pose prior.
-3 Our Proposed Method – Pose Prior Learner (PPL)
+Each memory bank explicitly contains multiple vectors encoding the variations of each sub-structure of a pose.
+This facilitates the aggregation of poses into a representative categorical pose prior.
+3 Our Proposed Pose Prior Learner (PPL)
 We introduce our proposed method, Pose Prior Learner (PPL).
-Given videos with static backgrounds featuring object instances from a specific category, such as dogs or humans, PPL can accurately estimate the poses of the objects in that category while gradually learning a general pose prior through unsupervised learning.
+Given images featuring object instances from a specific category, such as dogs or humans, PPL can accurately estimate the poses of the objects in that category while gradually learning a general pose prior through unsupervised learning.
 Note that our PPL requires no extra knowledge from human annotators.
@@
 Mathematically, we represent the topology for an object as a graph connecting keypoints with shared link weights, also known as connectivity.
-For any category, its general pose prior $V$ is defined as $V=(T,W)$, where $T$ represents its keypoint prior and $W$ denotes the connectivity prior.
-Specifically, $T$ consists of $N$ keypoints: $T=[P_{1},P_{2},...,P_{N}]$, where $P_{i}\in[-1,1]\times[-1,1]$ is the normalized pixel coordinates.
+For a category of objects, its general pose prior $V$ is defined as $V=(T,W)$, where $T$ represents its keypoint prior and $W$ denotes the connectivity prior.
+Specifically, $T$ consists of $N$ keypoints: $T=[P_{1},P_{2},...,P_{N}]$, where $P_{i}\in[-1,1]\times[-1,1]$ is the normalized 2D coordinates in the image pixel space.
+Unlike pre-defined priors, before training, $P_{i}$ does not explicitly encode the semantic parts of an object.
 $W$ is a $2D$ matrix of size $N\times N$, where each entry $w_{ij}$ in the matrix represents the connectivity probability between two keypoints $P_{i}$ and $P_{j}$.
 For instance, in the case of humans, the hand is connected to the torso via an arm; thus, the connectivity probability between these two parts should be higher than the connectivity probability between a hand and a foot.
-We randomly initialize the connectivity prior $W$ with positive values.
-The keypoint prior $T$ is decoded from a hierarchical memory $M$ storing compositional parts of prototypical poses.
+We initialize $W$ with random positive values.
+$T$ is decoded from a hierarchical memory $M$ storing compositional parts of prototypical poses, which is also randomly initialized.
 Figure 2: Overview of our proposed Pose Prior Learner (PPL).
 We first distill the keypoint prior from the hierarchical memory $M$.
-Features of the frame $I$ and the embedding of the keypoint prior are concatenated to predict the affine transformation parameters.
+Features of the image $I$ and the embedding of the keypoint prior are concatenated to predict the affine transformation parameters.
 The keypoint prior is transformed and their pair-wise links are modulated with the connectivity prior $W$ to obtain the combined link heatmap $S$.
-The concatenation of the link heatmap $S$ and the reference frame $I_{ref}$ is decoded to produce the reconstructed frame $I_{recon}$.
+The concatenation of the link heatmap $S$ and the reference image $I_{ref}$ is decoded to produce the reconstructed image $I_{recon}$.
 The $sg$ symbol represents the stopping gradient operation.
@@
 See Section 3.4 for training details.
-During training, PPL inputs a pair of frames $I$ and $I_{ref}$ with the same sizes of $\mathsf{H}\times\mathsf{W}\times 3$ where $\mathsf{H}$ and $\mathsf{W}$ are the height and the width of each video frame.
-The reference frame $I_{ref}$ is a randomly selected frame from the same video, containing the same person as $I$ but in a different pose.
-The aim of PPL is to learn to correctly predict the keypoints $T^{\prime}$ and their connectivity on $I$.
-Ideally, if PPL makes perfect predictions of the pose on $I$, by combining it with the background information from $I_{\text{ref}}$, the reconstructed image $I_{\text{recon}}$ should match $I$ exactly.
+During training, PPL inputs an image $I$ of $\mathsf{H}\times\mathsf{W}\times 3$ where $\mathsf{H}$ and $\mathsf{W}$ are the height and the width.
+The aim of PPL is to learn to correctly predict the image-specific keypoints $T^{\prime}$ and their connectivity on $I$ from the general categorical pose prior $V=(T,W)$.
+Ideally, if PPL makes perfect predictions of the pose on $I$, by combining it with the background information, the reconstructed image $I_{\text{recon}}$ should match $I$ exactly.
+The background information is provided by the reference image $I_{ref}$.
+For images extracted from video datasets, $I_{ref}$ can be a randomly selected frame from the same video that features the same object in a different pose.
+Alternatively, for a static image dataset, $I_{ref}$ can be a randomly masked version of the original image.
 Next, we introduce how we estimate $T^{\prime}$ on $I$ using the keypoint prior $T$.
 3.1 Transformation of the Keypoint Prior
-Given the frame $I$, we use a feature extractor $\phi_{enc}$ to extract its embedding $h_{I}$: $h_{I}=\phi_{enc}(I)$.
+Given the image $I$, we use a feature extractor $\phi_{enc}$ to extract its embedding $h_{I}$: $h_{I}=\phi_{enc}(I)$.
 $\phi_{enc}$ is a 2D-Convolution Neural Network (2D-CNN) trained from scratch.
@@
 Together with $h_{T}$ as inputs, PPL learns to predict the affine transformation parameters $\Theta_{i}\in[\Theta_{1},\Theta_{2},...,\Theta_{N}]$ for each keypoint $P_{i}$ in $T$ from $h_{I}$ via a two-layer fully connected network denoted as $FC(\cdot)$:
-$[\Theta_{1},\Theta_{2},...,\Theta_{i},...,\Theta_{N}]=FC(h_{I},h_{T}),\text{ where }\Theta_{i}=\begin{bmatrix}a^{(i)}&b^{(i)}&t_{x}^{(i)}\\ c^{(i)}&d^{(i)}&t_{y}^{(i)}\\ 0&0&1\end{bmatrix},$ (1)
-where $t_{x}^{(i)}$ and $t_{y}^{(i)}$ are the translations and $a^{(i)},b^{(i)},c^{(i)},d^{(i)}$ are the coefficients that define rotation, scaling, and shear.
-Each point $P_{i}$ in $T$ can then be transformed by $\Theta_{i}$, resulting in the keypoints $T^{\prime}$ for the frame $I$:
+$[\Theta_{1},\Theta_{2},...,\Theta_{i},...,\Theta_{N}]=FC(h_{I},h_{T}),$
+where $\text{where }\Theta_{i}=\begin{bmatrix}a^{(i)}&b^{(i)}&t_{x}^{(i)}\\ c^{(i)}&d^{(i)}&t_{y}^{(i)}\\ 0&0&1\end{bmatrix}.$ (1)
+$t_{x}^{(i)}$ and $t_{y}^{(i)}$ are the translations and $a^{(i)},b^{(i)},c^{(i)},d^{(i)}$ are the coefficients that define rotation, scaling, and shear.
+Each point $P_{i}$ in $T$ can then be transformed by $\Theta_{i}$, resulting in the image-specific keypoints $T^{\prime}$ for the image $I$:
 $T^{\prime}=[P^{\prime}_{1},P^{\prime}_{2},...,P^{\prime}_{N}],\text{ where }[P^{\prime}_{i},1]^{\top}=\Theta_{i}[P_{i},1]^{\top}.$ (2)
 3.2 Connecting Keypoints Based on the Connectivity Prior
-The connectivity of keypoints in objects is often fixed and rigid; for example, human arms maintain a relatively constant length, with a hand always connected to the torso via an arm and never connected to a foot.
+The connectivity of keypoints in objects is often fixed and rigid, for example, human arms maintain a relatively constant length, with a hand always connected to the torso via an arm and never connected to a foot.
 This rigidity in connectivity serves as a constraint, aiding in the regularization of pose estimation.
@@
 Ideally, if PPL correctly estimates the probability of physical links for an object category, $S_{i,j}$ will receive higher connectivity values, thereby activating the locations linking these two keypoints on the combined link map $S$.
-Given the combined link map $S$ and the reference frame $I_{ref}$, PPL can reconstruct the image $I$.
-$I_{ref}$ shares the same background as $I$ but features a different foreground pose, thereby offering background information for reconstruction.
-Meanwhile, $S$ provides the connectivity details among all the estimated keypoints.
-This encourages PPL to learn accurate pose estimations while minimizing the influence of background noise.
+Given the combined link map $S$ and the reference image $I_{ref}$, PPL can reconstruct the image $I$.
+$I_{ref}$ provides background information for reconstruction, while $S$ supplies foreground structural information by linking all the estimated keypoints with the connectivity prior.
 Therefore, we concatenate $S$ and $I_{ref}$ and feed them into a 2D-CNN to perform the image reconstruction $I_{recon}$, where $I_{recon}=\phi_{dec}(I_{ref},S)$.
 3.3 Reconstruction of Keypoint Configuration with Memory
-$M$ is a hierarchical memory storing compositional representations of prototypical poses.
-It contains $m$ memory banks $\{b_{1},b_{2},...,b_{m}\}$ of $k$ learnable tokens with $d$ randomly initialized parameters per token.
-$M$ is important to learn robust prior representations $V$ due to the following reasons: (1) By storing multiple prototypical poses, memory helps in aggregating these poses, which aids in creating a more robust and comprehensive prior that captures variations within an object category.
+$M$ is a memory module storing compositional representations of prototypical poses.
+$M$ is hierarchical because it consists of $m$ memory banks $\{b_{1},b_{2},...,b_{m}\}$, each containing $k$ learnable vectors of dimension $d$.
+This setup allows for efficient retrieval and management of information at multiple levels, where $M$ can represent pose configurations while memory banks capture specific pose sub-structures.
+This hierarchy enables $M$ to effectively learn robust keypoint prior representations $T$ for the following reasons: (1) By storing multiple prototypical poses, memory helps in aggregating these poses, which aids in creating a more robust and comprehensive prior that captures variations within an object category.
 (2) $M$ can assist in organizing information hierarchically, making it easier for PPL to retrieve relevant prototypical poses when making predictions in uncertain or ambugious scenarios, such as occlusion.
@@
 Next, we introduce how $M$ is structured.
-Given the estimated $N$ keypoints in $T^{\prime}$, we first encode it into $m$ tokens of dimension $d$ with several MLP-Mixer blocks (Tolstikhin et al., 2021), denoted as $MIX_{enc}$.
-We define $G$ as the collection of the $m$ tokens, where each token $g_{i}$ represents the embedding of a subset of the keypoints.
-Each token $g_{i}$ shares the same embedding space with the memory bank $b_{i}$ in memory $M$:
+Given the estimated $N$ keypoints in $T^{\prime}$, we first encode them into several tokens of dimension $d$ with several MLP-Mixer blocks (Tolstikhin et al., 2021), denoted as $MIX_{enc}$.
+The number of encoded tokens should always be equal to the number of memory banks, as we design these tokens to have different embedding space.
+Each token $g_{i}$ corresponds to the embedding of memory bank $b_{i}$ in memory $M$.
+We define $G$ as the collection of the $m$ tokens:
 $G=[g_{1},g_{2},...,g_{m}]=MIX_{enc}(T^{\prime}),$ (4)
-If all the tokens in all the memory banks of $M$ learn to capture unique parts of prototypical poses, each $g_{i}$ should be able to retrieve the most similar token in the memory bank $b_{i}$ of $M$ and reconstructs $G$ itself.
-Here, we use $L2$ distance to measure the similarity between $g_{i}$ and $k$ token in $b_{i}$ of $M$.
-The token in $b_{i}$ that is most similar to $g_{i}$ is denoted as $g^{\prime}_{i}$.
+If all the vectors in all the memory banks of $M$ learn to capture unique parts of prototypical poses, each $g_{i}$ should be able to retrieve the most similar vector in the memory bank $b_{i}$ of $M$ and reconstruct $G$ itself.
+Here, we use $L2$ distance to measure the similarity between $g_{i}$ and $k$ vectors in $b_{i}$ of $M$.
+The vector in $b_{i}$ that is most similar to $g_{i}$ is denoted as $g^{\prime}_{i}$.
 Thus, for each $g_{i}$, we can always find the corresponding $g^{\prime}_{i}$ from $b_{i}$, resulting in a collection of $G^{\prime}=[g^{\prime}_{1},g^{\prime}_{2},...,g^{\prime}_{m}]$.
 Another series of MLP-Mixer blocks (denoted as $MIX_{dec}$) is then used to decode $G^{\prime}$ back to $N$ keypoints $T^{\prime}_{recon}$, where $T^{\prime}_{recon}=MIX_{dec}(G^{\prime})$.
-See Figure LABEL:fig:mem_stru for the schematic.
-Unlike PCT (Geng et al., 2023) where all tokens of a memory are within the same embedding space, We organize $M$ into $m$ memory banks, each containing $k$ tokens that represent an independent embedding space.
-This structure enables us to efficiently distill compositional parts from each memory bank, allowing us to form a general representation of a pose, which is distilled into the keypoint prior $T$.
+See Figure ALABEL:fig:mem_stru in Appendix A5 for the schematic.
+Unlike PCT (Geng et al., 2023) where all vectors of a memory are within the same embedding space, We organize $M$ into $m$ memory banks, each containing $k$ vectors that represent an independent embedding space.
+This structure enables us to efficiently distill each memory bank to form a general pose representation, which can be viewed as the keypoint prior $T$.
 Specifically, we define $MP(\cdot)$ as the mean pooling operation.
-PPL pools one token vector among the $k$ vectors of each memory bank.
-These pooled $m$ vectors are further decoded by $MIX_{dec}$ into $N$ points of our distilled keypoint prior $T$.
-See Figure LABEL:fig:mem_cons for the schematic of distilling $T$: $T=MIX_{dec}([MP(b_{1}),MP(b_{2}),...,MP(b_{m})])$.
+PPL pools the $k$ vectors of each memory bank into one vector.
+These pooled $m$ vectors are further decoded by $MIX_{dec}$ into $N$ points of our keypoint prior $T$.
+See Figure ALABEL:fig:mem_cons in Appendix A5 for the schematic of distilling $T$: $T=MIX_{dec}([MP(b_{1}),MP(b_{2}),...,MP(b_{m})])$.
+Compared to simply using a moving average of the learned poses, the advantages of employing hierarchical memory and keypoint configuration reconstruction are as follows: (1) the hierarchical memory increases the complexity of the embedding space with compositional parts, enhances the expressive capability for poses and facilitates more accurate pose retrieval for occluded images; (2) the compositional nature of the hierarchical memory addresses data bias, enabling the distillation of a less biased prior.
 3.4 Training and Inference
@@
 We elaborate on these four losses below.
+Additionally, we use three training techniques for stable convergence of the network.
+We present the details of these training techniques in Appendix A6.
 Image Reconstruction Loss.
@@
 To ensure that the network does not transform the points in the keypoint prior outside the boundaries of the image, we limit the x and y coordinates of the transformed keypoints to be within the image:
-$L_{b}=\sum_{*\in{x,y}}\begin{cases}|P^{\prime}_{i,*}|&\text{if }|P^{\prime}_{i,*}|>1,\\ 0&\text{otherwise}.\end{cases}\text{ and }L_{l}=\sum_{i,j}w_{i,j}\|l(P_{i},P_{j})-l(P^{\prime}_{i},P^{\prime}_{j})\|_{1},$ (5)
+$L_{b}=\sum_{*\in{x,y}}\begin{cases}|P^{\prime}_{i,*}|&\text{if }|P^{\prime}_{i,*}|>1,\\ 0&\text{otherwise}.\end{cases}$ (5)
 where $P^{\prime}_{i,x}$ and $P^{\prime}_{i,y}$ are the normalized $x$ and $y$ coordinate of the keypoint $P^{\prime}_{i}$ respectively.
@@
 Thus, we propose the constraint that links should be assigned a high weight if they do not vary significantly in length before and after the affine transformation.
-The loss $L_{l}$ encourages the preservation of link lengths during pose estimation.
-It is defined as in the equation above, where $l(\cdot)$ is the $L2$ distance between two keypoints before and after the affine transformation.
+The loss $L_{l}$ encourages the preservation of the link lengths during pose estimation.
+It is defined as in the equation below, where $l(\cdot)$ is the $L2$ distance between two keypoints before and after the affine transformation:
+$L_{l}=\sum_{i,j}w_{i,j}\|l(P_{i},P_{j})-l(P^{\prime}_{i},P^{\prime}_{j})\|_{1}.$ (6)
 Reconstruction Loss on Keypoint Configurations.
-In Section 3.3, given a collection of token representations in $G$, PPL retrieves the most similar tokens from each memory bank of the hierarchical memory $M$ and generates $G^{\prime}$ in a non-differentiable manner.
-To ensure that $M$ learns to store meaningful token embeddings that represent compositional parts of poses, the retrieved token representations $G^{\prime}$ from $M$ should closely match $G$.
-Moreover, if these compositional parts are structured correctly, the tokens should be able to decode into meaningful keypoint configurations $T^{\prime}_{recon}$ that are close to the original keypoint configurations $T^{\prime}$.
-Therefore, we introduce the keypoint configuration reconstruction loss $L{kr}$, defined as: $L_{kr}=\|T^{\prime}_{recon}-T^{\prime}\|_{2}+\|G-G^{\prime}\|_{2}$.
-Training Techniques.
-To ensure convergence and stability during the training of PPL, we introduce three gradient dettachment techniques: (1) To address the broken gradient issue during the quantization step from $G$ to $G^{\prime}$, we adopt the approach from VQ-VAE (Van Den Oord et al., 2017).
-Specifically, our PPL copies the gradients of $G^{\prime}$ to $G$ for backward propagation, allowing the gradients to flow through the quantization step.
-(2) The hierarchical memory $M$ is updated using an exponential moving average to smooth the gradient updates, particularly during the early stages of training when $G$ can be quite noisy.
-This approach helps stabilize the learning process and ensures that $M$ retains more reliable information over time.
-(3) For $M$ to learn effective representations of $G$, it requires an accurate estimation of $T^{\prime}$, which depends on a good prior $V$ distilled from $M$.
-This creates a chicken-and-egg problem that complicates training.
-To address this, we introduce two gradient detachments to separate the training processes.
-First, we detach the gradients from $T$ and train the keypoint transformation and image reconstruction pathway, as shown by the red arrows in Figure 2.
-Second, we detach the gradients from $T^{\prime}$ to train the memory encoder and decoder, $MIX_{\text{enc}}$ and $MIX_{\text{dec}}$, as indicated by the green arrows in Figure LABEL:fig:mem_stru.
-Iterative Inference.
-We propose an iterative inference strategy (Figure 4).
-$4$ iterations are used for every experiment.
-In every iteration, we take the reconstructed frame $I_{recon}$ from the last iteration (the original frame $I$ for iteration 00) as the input.
-We infer its keypoints $T^{\prime}$ as the output keypoints of the current iteration.
-The hierarchical memory $M$ is used to reconstruct $T^{\prime}$ and the reconstructed keypoints $T^{\prime}_{recon}$ are used to obtain the reconstructed frame $I_{recon}$.
-$I_{recon}$ is then used as the input for the next iteration.
-We keep the original occluded frame $I$ as the reference frame for all the iterations.
-See Appendix A1 for more implementation details.
-Figure 4: Overview of the iterative inference strategy in our PPL.
-During inference, we iteratively use the reconstructed frame $I_{recon}$ as input to estimate the pose $T^{\prime}$.
-The hierarchical memory $M$ refines the estimated pose $T^{\prime}$ and outputs $T^{\prime}_{recon}$.
-With the original test image $I$ as the reference frame $I_{ref}$, PPL reconstructs the image $I_{recon}$.
-It is then used as the input frame in the next iteration.
-See Section 3.4 for details.
-4 Experiments
-We use three video datasets: Human3.6m (Ionescu et al., 2013), Taichi (Siarohin et al., 2019), and YouTube dog videos with green backgrounds for our experiments.
-See Appendix A2 for details.
-On Human3.6m, we report the results in the mean $L2$ error between the predicted keypoints and the ground truth, normalized by the image size.
-For Taichi, as in (He et al., 2022a; Siarohin et al., 2021; He et al., 2021; 2022b; Zhang et al., 2018) we use the summed $L2$ error computed at a resolution of $256\times 256$.
-To ensure a more accurate evaluation, unlike previous methods (He et al., 2022a; Schmidtke et al., 2021), we employ a minimum flow algorithm (Waissi, 1994) to align the predicted keypoints with the ground truth.
+In Section 3.3, given a collection of token representations in $G$, PPL retrieves the most similar vectors from each memory bank of the hierarchical memory $M$ and generates $G^{\prime}$ in a non-differentiable manner.
+To ensure that $M$ learns to store meaningful token embeddings that represent compositional parts of poses, the retrieved vectors $G^{\prime}$ from $M$ should closely match $G$.
+Moreover, if these compositional parts are structured correctly, the vectors should be able to decode into meaningful keypoint configurations $T^{\prime}_{recon}$ that are close to the original keypoint configurations $T^{\prime}$.
+Therefore, we introduce the keypoint configuration reconstruction loss defined as: $L_{kr}=\|T^{\prime}_{recon}-T^{\prime}\|_{2}+\|G-G^{\prime}\|_{2}$.
+Table 1: Keypoint detection on CUB-200-2011.
+We report the mean $L2$ error normalized by the image resolutions of $128\times 128$.
+We use 10 keypoints for CUB-Aligned and 4 keypoints for CUB-001, CUB-002, CUB-003, and CUB-all.
+The best is in bold.
+Method CUB- aligned CUB- 001 CUB- 002 CUB- 003 CUB- all
+(Zhang et al., 2018) 5.36 26.9 27.6 27.1 22.4
+(He et al., 2021) 5.21 22.6 29.1 21.2 14.7
+(He et al., 2022b) 3.23 22.1 22.3 21.5 12.1
+AutoLink (He et al., 2022a) 3.51 20.2 19.2 18.5 11.3
+PPL (ours) 3.19 19.3 18.6 17.3 10.5
 Table 2: Keypoint detection on Human3.6m.
 We report the mean $L2$ error normalized by the image resolutions (Res.) of both $256\times 256$ and $128\times 128$.
-* is the $L2$ error calculated by our minimum flow algorithm.
+* is the result trained by using masked images as reference images.
 The best is in bold.
@@
 AutoLink (He et al., 2022a) 128 2.76
-AutoLink* (He et al., 2022a) 128 2.61
 PPL (ours) 128 1.92
+PPL* (ours) 128 2.23
 PPL (ours) 256 2.56
@@
 For consistency with the baseline methods, we report the summed $L2$ error at a resolution of $256\times 256$.
-* is the $L2$ error calculated by our minimum flow algorithm.
+* is the result trained by using masked images as reference images.
 The best is in bold.
@@
 AutoLink (He et al., 2022a) 316.10
-AutoLink* (He et al., 2022a) 302.41
+PPL* (ours) 298.60
 PPL (ours) 293.35
-4.1 Unsupervised Human Pose Estimation
-We compare the keypoint detection results of PPL with other unsupervised pose estimation methods and present the results in Table 1 (Human3.6m) and Table 2 (Taichi).
-On both datasets, PPL significantly outperforms all baselines across all image resolutions.
-Among the baselines, AutoLink (He et al., 2022a) also incorporates learnable connectivity priors, similar to our PPL.
+Iterative Inference.
+We propose an iterative inference strategy (Figure 3).
+$4$ iterations are used for every experiment.
+In every iteration, we take the reconstructed image $I_{recon}$ from the last iteration (the original image $I$ for iteration 00) as the input.
+We infer its keypoints $T^{\prime}$ as the output keypoints of the current iteration.
+The hierarchical memory $M$ is used to reconstruct $T^{\prime}$ and the reconstructed keypoints $T^{\prime}_{recon}$ are used to obtain the reconstructed image $I_{recon}$.
+$I_{recon}$ is then used as the input for the next iteration.
+We keep the original occluded image $I$ as the reference image for all the iterations.
+See Appendix A1 for more implementation details.
+Figure 3: Overview of the iterative inference strategy in our PPL (Section 3.4).
+During inference, we iteratively use the reconstructed image $I_{recon}$ as input to estimate the pose $T^{\prime}$.
+The hierarchical memory $M$ refines the estimated pose $T^{\prime}$ and outputs $T^{\prime}_{recon}$.
+With the original image $I$ as the reference image $I_{ref}$, PPL reconstructs the image $I_{recon}$.
+It is then used as the input image in the next iteration.
+4 Experiments
+For quantitative experiments, we use three image datasets: Human3.6m (Ionescu et al., 2013), Taichi (Siarohin et al., 2019b, a), and CUB-200-2011 (Wah et al., 2011).
+For fair comparisons, we set the number of keypoints $N$ to be the same as (He et al., 2022a).
+For qualitative visualizations, we use Youtube dog videos, Flowers (Nilsback & Zisserman, 2008), 11k-Hands (Afifi, 2019), and Horses (Zhu et al., 2017).
+See Appendix A2 for details.
+We utilize video frames as $I_{ref}$ for the Human3.6m, Taichi, and YouTube dog videos, while randomly masked images are used as $I_{ref}$ for other datasets.
+On Human3.6m and CUB-200-2011, we report the results in the mean $L2$ error between the predicted keypoints and the ground truth, normalized by the image size.
+For Taichi (He et al., 2022a; Siarohin et al., 2021; He et al., 2021, 2022b; Zhang et al., 2018), we use the summed $L2$ error computed at a resolution of $256\times 256$.
+4.1 Unsupervised Categorical Pose Estimation
+We compare the keypoint detection results of PPL with other unsupervised pose estimation methods and present the results in Table 1 (CUB-200-2011), Table 2 (Human3.6m), and Table 3 (Taichi).
+On all datasets, PPL significantly outperforms all baselines across all image resolutions.
+Among the baselines, AutoLink (He et al., 2022a) also incorporates learnable connectivity priors.
 However, its performance is inferior due to the absence of hierarchical memory.
-Additionally, we note that STT (Schmidtke et al., 2021), a baseline with human-defined priors, still underperforms compared to our PPL.
-Consistent with HPE (Yoo & Russakovsky, 2023), this suggests that pre-defined priors are not always optimal, and our PPL is able to learn more representative priors that outperform those manually defined.
+Notably, even when masked images are used as $I_{ref}$, PPL consistently outperforms AutoLink.
+This demonstrates that the effectiveness of PPL is primarily attributed to its pose prior, rather than relying on cross-frame reconstruction.
+Additionally, we note that STT (Schmidtke et al., 2021), a baseline with human-defined priors, still underperforms PPL.
+Consistent with HPE (Yoo & Russakovsky, 2023), this suggests that pre-defined priors are not always optimal.
+PPL can learn more representative priors that outperform those manually defined.
 Visualization of Pose Estimation.
-We provide the visualization results of pose estimation on Human3.6m (Figure LABEL:fig:human-pose-vis) and YouTube dog videos (Figure LABEL:fig:dog-pose-vis).
-The visualization results demonstrate that PPL can learn priors and estimate poses of various object categories, without any external knowledge.
-For example, in Row 3, Column 2 of Figure LABEL:fig:human-pose-vis, PPL correctly estimates the bowing pose of a person.
-In Row 3, Column 5 of Figure LABEL:fig:dog-pose-vis, PPL correctly estimates the pose of a dog lowering down its head.
+We provide the visualization results of pose estimation on Human3.6m (Figure ALABEL:human-vis), YouTube dog videos (Figure ALABEL:dog-vis), 11k-Hands (Figure ALABEL:hand-vis), CUB-200-2011 (Figure ALABEL:bird-vis), Horse (Figure ALABEL:horse-vis), and Flowers (Figure ALABEL:flower-vis) in Appendix A7.
+The visualization results demonstrate that PPL can learn categorical priors and estimate poses for various object categories, without any external annotations.
+For example, in Row 2, Column 2 of Figure ALABEL:human-vis, PPL correctly estimates the bowing pose of a person.
+In Row 2, Column 5 of Figure ALABEL:dog-vis, PPL correctly estimates the pose of a dog lowering down its head.
 Moreover, we found the quality of estimated poses of dogs is inferior to that of humans.
@@
 Additionally, dogs often perform actions such as turning around, which can lead to significant changes in pose that are difficult to accurately capture by our priors in 2D space.
+Additional visualization results for Flowers, Hands, and Horses in Appendix A7 further demonstrate that our PPL can be applied for various categories.
+Notably, flowers are rigid objects with no degrees of freedom like animals.
+Yet, our PPL still learns meaningful categorical priors for them.
 Visualization of the Pose Prior Changing with the Training Epochs.
@@
 Notably, the learnable keypoints align with the human joints defined in the literature, and the connectivity among keypoints corresponds to the physical connections between body parts.
-As training continues, the connectivity prior gradually learns the skeletal structure of the human body, with irrelevant links between keypoints diminishing over time (as seen when comparing epochs 15 and 20).
+As training continues, the connectivity prior gradually learns the skeletal structure of the human body, with irrelevant links between keypoints diminishing over time, as seen when comparing epochs 15 and 20.
 4.2 Ablation Studies
@@
 (5) In contrast to (4), freezing random connectivity priors prevents the model from converging, implying that connectivity priors play a more critical role in guiding pose estimations than keypoint priors.
-Ablation on Number & Dimension of Tokens in Each Memory Bank.
+Ablation on Number & Dimension of Vectors in Each Memory Bank.
 In our hierarchical memory, we used 34 memory banks.
-Here, we analyze the impact of the number of tokens per memory bank and the dimension of each token on PPL’s pose estimation performance.
-From Figure A1 in Appendix A3, we observed that PPL remains robust across different token counts and dimensions, although performance slightly improves with more tokens of higher dimensions.
-As a result, we fixed 16 tokens per memory bank, each with a dimension of 512, for all experiments.
+Here, we analyze the impact of the number of vectors per memory bank and the dimension of each vector on PPL’s pose estimation performance.
+From Figure 5 in Appendix A3, we observed that PPL remains robust across different vector counts and dimensions, although performance slightly improves with more vectors of higher dimensions.
+As a result, we fixed 16 vectors per memory bank, each with a dimension of 512, for all experiments.
 Ablation on Number of Keypoints.
 We varied the number of keypoints in the pose priors from 4 to 32.
-The results in Figure A1 in Appendix A3 show that pose estimation accuracy improves as the number of keypoints in the prior increases.
-However, using 32 keypoints offers limited improvement compared to PPL with 16 keypoints for human pose estimation.
-4.3 Pose Estimation in Occlusion Scene
+The results in Figure 5 in Appendix A3 show that pose estimation accuracy improves as the number of keypoints in the prior increases.
+However, using 32 keypoints offers limited improvement compared to PPL with 16 keypoints.
+4.3 Pose Estimation in Occluded Scenes
 To verify the robustness of PPL in occluded scenes, we divide the image into $32\times 32$ patches and apply two masking techniques: RandomMasking and CenterMasking.
@@
 We explore the effect of occluded areas on PPL.
-From Figure 9 and Figure 9, we observe that at iteration 0, as the occluded areas increase, overall performance declines with larger occlusions.
+From Figure A2 and Figure A2 in Appendix A4, we observe that at iteration 0, as the occluded areas increase, overall performance declines with larger occlusions.
 However, with our iterative inference strategy, PPL effectively infers the missing parts of the poses by utilizing prototypical poses stored in hierarchical memory and the learned priors.
 Notably, it restores partially occluded poses to reasonably complete full-body poses, leading to a lower $L2$ error, comparable to those without occlusion.
-This effect is more pronounced when the occlusion area is small or medium.
+This effect is more pronounced with smaller occluded areas.
 Visualization of Pose Estimation with Occlusion.
 We present the estimated poses by our PPL for occluded images as a function of the number of inference iterations in Figure LABEL:fig:occlusion-vis.
-Across both RandomMasking (top 2 rows) and CenterMasking (bottom 2 rows), with our iterative inference strategy, PPL successfully reconstructs the occluded part of the image after three iterations and meanwhile, predicts reasonable complete full-body poses.
-Figure 8: PPL Results of Keypoint Detection as a Function of Number of Inference Iterations on Images with RandomMasking from Human3.6m.
-The “mask ratio” in the legend specifies the masked proportion on the $32\times 32=1024$ image patches.
-Figure 9: PPL Results of Keypoint Detection as a Function of Number of Inference Iterations on Images with CenterMasking from Human3.6m.
-The “mask size” in the legend refers to the width and height of the masked region, on $1024$ image patches.
+Across both RandomMasking and CenterMasking, with our iterative inference strategy, PPL successfully reconstructs the occluded image parts after three iterations and meanwhile, predicts reasonable full-body poses.
 5 Discussion
-We introduce the challenge of unsupervised prior learning and highlight its significance in pose estimation.
+We introduce the challenge of unsupervised categorical prior learning and highlight its significance in pose estimation.
 To address this, we propose a novel method called Pose Prior Learner (PPL).
 PPL utilizes a hierarchical memory to store compositional parts of learnable prototypical poses, which are distilled into a general pose prior for any object category.
-Our experimental results show that PPL requires no additional knowledge and outperforms recent competitive baselines in video pose estimation.
+Our experimental results show that PPL requires no additional human annotations and outperforms recent competitive baselines in pose estimation.
 Notably, the learned prior proves to be even more effective in pose estimation than methods that rely on human-defined priors.
 With hierarchical memory and learned priors, PPL can perform iterative inferences and robustly estimate poses in occluded scenes.
-Despite outstanding performance in unsupervised pose estimation, PPL has several limitations.
-For instance, it requires video data captured in static backgrounds.
-Additionally, PPL learns 2D priors, which makes it difficult to capture real-world 3D postures.
-Thus, PPL struggles in scenarios where foreground objects involve rotations or significant shape changes.
+Despite outstanding performance in unsupervised categorical pose estimation, PPL has several limitations.
+For instance, it learns 2D priors, which makes it difficult to capture real-world 3D postures.
+Thus, PPL struggles in scenarios where objects involve rotations or significant shape changes.
 Extending PPL to incorporate 3D priors will be a key focus of our future research.
+6 Impact Statement
+We present the challenge of unsupervised learning of categorical priors for pose estimation and propose Pose Prior Learner (PPL) to address it.
+By leveraging learned priors, PPL enhances robustness and generalizability during inference, promoting the safe deployment of AI technologies.
+However, privacy concerns surrounding images and videos in these datasets, especially for human pose estimation, must be carefully addressed.
+Protecting personal and sensitive information remains paramount as AI advances.
 Appendix A1 Implementation Details of our Pose Prior Learner
@@
 We conduct experiments using the link thickness $\sigma^{2}=5\times 10^{-4}$ across all benchmark datasets, where we adopt the same definition of $\sigma$ used in (He et al., 2022a).
-For the hierarchical memory, we use $34$ memory banks, each of which contains $16$ tokens of dimension $512$, for all experiments.
+For the hierarchical memory, we use $34$ memory banks, each of which contains $16$ vectors of dimension $512$, for all experiments.
 Appendix A2 Datasets
@@
 For training, we use subjects 1, 5, 6, 7, 8, and 9, and for testing, we use subject 11.
-Taichi (Siarohin et al., 2019) consists of 3,049 training videos and 285 test videos featuring individuals performing Tai-Chi, with diverse foreground and background appearances.
+Taichi (Siarohin et al., 2019b, a) consists of 3,049 training videos and 285 test videos featuring individuals performing Tai-Chi, with diverse foreground and background appearances.
 Following the approach in (Siarohin et al., 2021), we use 5,000 frames for training and 300 frames for testing.
-YouTube dog videos are videos with green backgrounds collected from YouTube to further qualitatively demonstrate the performance of PPL in non-human species.
-We used only 20 videos from YouTube and decomposed them into a total of 2000 video frames for training.
-We use this customized dataset to demonstrate the applicability of PPL in learning pose priors for non-human categories.
+YouTube dog videos are videos with green backgrounds collected from YouTube to further qualitatively demonstrate the performance of PPL.
+Existing dog datasets are not suitable, as they often contain multiple, partially occluded dog instances.
+Therefore, we curated a custom dataset using 20 YouTube videos, extracting 2,000 frames for training.
+This dataset allows us to demonstrate PPL’s ability to learn pose priors for non-human categories without using ground truth poses.
 All images are trained and tested at a resolution of $256\times 256$.
 We use $10$ keypoints for this category and provide the visualization of the estimated poses on the test videos.
-Appendix A3 Ablation on Memory Bank Tokens and Number of Keypoints
-Figure A1: Ablation of our PPL method on Memory Bank Tokens and Number of Keypoints in Human3.6m.
-The upper horizontal axis is the number of keypoints (ranging from 4 to 32) and the lower horizontal axis is the dimension of memory bank tokens (ranging from 64 to 512).
-The dashed purple line is for ablations on number of keypoints and the solid lines are for ablations on memory bank tokens.
+The dataset will be publicly released together with other data, models, and source code.
+CUB-200-2011 (Wah et al., 2011) contains 11788 images of birds.
+We crop and align the images according to (He et al., 2022a).
+We use the train/val/test split of (Choudhury et al., 2021).
+All images are trained and tested at a resolution of $128\times 128$.
+11k-Hands (Afifi, 2019), Horse (Zhu et al., 2017), and Flowers (Nilsback & Zisserman, 2008) are used for qualitative visualization.
+Images with multiple horses are removed and all horses are aligned to face left.
+All images are trained and tested at a resolution of $128\times 128$.
+Appendix A3 Ablation on Memory Bank Vectors and Keypoint Numbers
+Figure A0: Ablation of our PPL method on memory bank vectors and number of keypoints in Human3.6m.
+The upper horizontal axis is the number of keypoints (ranging from 4 to 32) and the lower horizontal axis is the dimension of memory bank vectors (ranging from 64 to 512).
+The dashed purple line is for ablations on number of keypoints and the solid lines are for ablations on memory bank vectors.
+Appendix A4 PPL Results of Keypoint Detection on Occluded Images
+Figure A1: PPL results of keypoint detection as a function of number of inference iterations on images with RandomMasking from Human3.6m.
+The “mask ratio” in the legend specifies the masked proportion on the $32\times 32=1024$ image patches.
+Figure A2: PPL results of keypoint detection as a function of number of inference iterations on images with CenterMasking from Human3.6m.
+The “mask size” in the legend refers to the width and height of the masked region, on $1024$ image patches.
+Appendix A5 Retrieval and Distillation of the proposed Hierarchical Memory in PPL
+Appendix A6 Training Techniques
+To ensure convergence and stability during the training of PPL, we introduce three gradient dettachment techniques: (1) To address the broken gradient issue during the quantization step from $G$ to $G^{\prime}$, we adopt the approach from VQ-VAE (Van Den Oord et al., 2017).
+Specifically, our PPL copies the gradients of $G^{\prime}$ to $G$ for backward propagation, allowing the gradients to flow through the quantization step.
+(2) The hierarchical memory $M$ is updated using an exponential moving average to smooth the gradient updates, particularly during the early stages of training when $G$ can be quite noisy.
+This approach helps stabilize the learning process and ensures that $M$ retains more reliable information over time.
+(3) For $M$ to learn effective representations of $G$, it requires an accurate estimation of $T^{\prime}$, which depends on a good prior $V$ distilled from $M$.
+This creates a chicken-and-egg problem that complicates training.
+To address this, we introduce two gradient detachments to separate the training processes.
+First, we detach the gradients from $T$ and train the keypoint transformation and image reconstruction pathway, as shown by the red arrows in Figure 2.
+Second, we detach the gradients from $T^{\prime}$ to train the memory encoder and decoder, $MIX_{\text{enc}}$ and $MIX_{\text{dec}}$, as indicated by the green arrows in Figure ALABEL:fig:mem_stru in Appendix A5.
+Appendix A7 Additional Visualizations

### v3

@@
 Priors represent beliefs or assumptions about a system or the characteristics of a concept.
-They are widely used in statistical inference (Lindley, 1961), cognitive science (Schad et al., 2021), and machine learning (Diligenti et al., 2017; Gülçehre & Bengio, 2016).
-This pre-existing knowledge is essential for guiding the inference process, enabling AI models to make robust predictions in uncertain or ambiguous situations (Thiruvenkadam et al., 2008; Sung et al., 2015; Liang et al., 2024).
+They are widely used in statistical inference Lindley (1961), cognitive science Schad et al.
+(2021), and machine learning Diligenti et al.
+(2017); Gülçehre & Bengio (2016).
+This pre-existing knowledge is essential for guiding the inference process, enabling AI models to make robust predictions in uncertain or ambiguous situations Thiruvenkadam et al.
+(2008); Sung et al.
+(2015); Liang et al.
+(2024).
 The objective of our work is to enhance our understanding of priors in AI models and offer preliminary answers to the three key intelligence questions: (1) How do we acquire priors in the first place?
@@
 It encapsulates the expectation of the keypoint configurations and the connectivity between keypoints.
-In parallel to our challenge of unsupervised categorical prior learning from images for pose estimation, unsupervised pose estimation leverages the abundant, unannotated visual information available in large image datasets to extract pose information (Hu & Ahuja, 2021; Sommer et al., 2024; Chen et al., 2019; He et al., 2022a; Schmidtke et al., 2021).
+In parallel to our challenge of unsupervised categorical prior learning from images for pose estimation, unsupervised pose estimation leverages the abundant, unannotated visual information available in large image datasets to extract pose information Hu & Ahuja (2021); Sommer et al.
+(2024b); Chen et al.
+(2019); He et al.
+(2022a); Schmidtke et al.
+(2021).
 The use of pose priors can provide valuable guidance in this process.
 We categorize the existing works in unsupervised pose estimation into two groups: those that incorporate hand-made priors and those that operate without any priors.
-Recent approaches (He et al., 2022a; Sun et al., 2022, 2023) attempt to predict keypoints from images, construct object structure representations using these keypoints, and learn effective structural information through image reconstruction.
+Recent approaches He et al.
+(2022a); Sun et al.
+(2022; 2023) attempt to predict keypoints from images, construct object structure representations using these keypoints, and learn effective structural information through image reconstruction.
 However, without pose priors, these methods can be disrupted by background information or may predict infeasible topological configurations of an object during occlusion.
 The risk of generating inaccurate keypoints stems from the absence of supplementary information that could help refine both keypoint localization and the connections between them.
-The other group of methods (Schmidtke et al., 2021; Yoo & Russakovsky, 2023) utilize prior knowledge of a category’s general pose to guide the pose estimation of individuals within that category.
+The other group of methods Schmidtke et al.
+(2021); Yoo & Russakovsky (2023) utilize prior knowledge of a category’s general pose to guide the pose estimation of individuals within that category.
 Conceptually, each category is expected to exhibit a generalized and distinctive pose prior that reflects characteristics such as shape, size, and structure.
@@
 As training progresses, the hierarchical memory retains and aggregates multiple accurate prototypical poses, thereby contributing to a more precise pose prior and enhancing the model’s ability to estimate poses.
+Similar to Wu et al.
+(2023) and Li et al.
+(2024), PPL learns object poses via image-level reconstruction, but uniquely does so from unannotated images without any object masks.
+Moreover, although methods such as Sommer et al.
+(2024a) also tackle category-level pose estimation, they rely on dense point clouds for scene reconstruction.
+In contrast, PPL leverages sparse keypoints and their relations, yielding a lightweight, semantic, and invariant pose representation well-suited for high-level scene understanding.
 Upon completing the training, we obtain a model that enables accurate pose estimation, a categorical pose prior that encapsulates the general features of a category, and a hierarchical memory that stores diverse prototypical poses for that category.
@@
 Our contributions are highlighted below:
-1. We introduce the challenge of unsupervised categorical prior learning in the context of pose estimation.
+1. We introduce the challenge of unsupervised categorical prior learning for pose estimation.
 2.
@@
 Experimental results demonstrate that our PPL accurately estimates poses, even in occluded scenes.
+Figure 2: Overview of our proposed Pose Prior Learner (PPL).
+We first distill the keypoint prior from the hierarchical memory $M$.
+Features of the image $I$ and the embedding of the keypoint prior are concatenated to predict the affine transformation parameters.
+The keypoint prior is transformed and their pair-wise links are modulated with the connectivity prior $W$ to obtain the combined link heatmap $S$.
+The concatenation of the link heatmap $S$ and the reference image $I_{ref}$ is decoded to produce the reconstructed image $I_{recon}$.
+The sgsg symbol represents the stopping gradient operation.
+The red arrows indicate the gradient flows during backpropagation based on image reconstruction.
+See Section 3.4 for training details.
 2 Related Works
 Unsupervised Pose Estimation without Priors.
-Numerous unsupervised learning methods without priors have been proposed to detect keypoints from images, which are then used to reconstruct images for supervision (Li et al., 2021; Geng et al., 2021; Zhang et al., 2018; Sun et al., 2022; Thewlis et al., 2017; Jakab et al., 2020).
-For example, AutoLink (He et al., 2022a) extracts keypoints from the image and estimates the strength of the links between pairs of keypoints.
+Numerous unsupervised learning methods without priors have been proposed to detect keypoints from images, which are then used to reconstruct images for supervision Li et al.
+(2021); Geng et al.
+(2021); Zhang et al.
+(2018); Sun et al.
+(2022); Thewlis et al.
+(2017); Jakab et al.
+(2020).
+For example, AutoLink He et al.
+(2022a) extracts keypoints from the image and estimates the strength of the links between pairs of keypoints.
 It then combines these keypoints with the link heatmap to reconstruct the randomly masked image.
 In these methods, keypoints are directly predicted from the image and supervised solely by image reconstruction, leading to potential detection of keypoints in background regions with complex textures.
-To alleviate this problem, BKind (Sun et al., 2022, 2023) uses keypoints extracted from two video frames to reconstruct the pixel-level differences between these two frames.
+To alleviate this problem, BKind Sun et al.
+(2022; 2023) uses keypoints extracted from two video frames to reconstruct the pixel-level differences between these two frames.
 However, the lack of constraints on keypoint configuration and connectivity still undermines the reliability of their approach.
@@
 Unsupervised Pose Estimation Incorporating Human-defined Priors.
-Several methods utilize prior knowledge from human annotators to guide the pose estimation (Chen & Dou, 2021; Shi et al., 2023; Zhang et al., 2022; Schmidtke et al., 2021; Yoo & Russakovsky, 2023).
-Among these methods, Shape Template Transforming (STT) (Schmidtke et al., 2021) applies affine transformations to a predefined pose prior, aligning it with the estimated pose from a video frame.
+Several methods utilize prior knowledge from human annotators to guide the pose estimation Chen & Dou (2021); Shi et al.
+(2023); Zhang et al.
+(2022); Schmidtke et al.
+(2021); Yoo & Russakovsky (2023).
+Among these methods, Shape Template Transforming (STT) Schmidtke et al.
+(2021) applies affine transformations to a predefined pose prior, aligning it with the estimated pose from a video frame.
 By incorporating an additional frame from the same video to provide background information, an image reconstruction loss supervises the pose estimation process.
@@
 However, pose priors are often difficult to obtain, requiring costly human annotations.
-Moreover, HPE (Yoo & Russakovsky, 2023) has shown that predefined pose priors are not always optimal, and tuning the shape of the prior can sometimes improve performance.
+Moreover, HPE Yoo & Russakovsky (2023) has shown that predefined pose priors are not always optimal, and tuning the shape of the prior can sometimes improve performance.
 Unlike these methods, our approach learns the prior directly from input images without any manual annotations, and models with our learned priors even outperform those using human-defined priors.
-Compositional Memory Architectures.
-Compositional memory has been widely used in many computer vision tasks, such as question answering (Seong et al., 2021), object segmentation (Seong et al., 2021), and sence graph generation (Deng et al., 2022).
-In pose estimation, PCT (Geng et al., 2023) decomposes a human pose into discrete tokens, where each token connects several interdependent joints and characterizes a sub-structure of the entire human pose.
-This approach is highly effective for decomposing and reconstructing poses, providing robust pose representations while significantly reducing computation and storage costs.
-However, PCT encodes all tokens into the same embedding space, making it difficult to aggregate semantic tokens that represent different sub-structures.
-In contrast, our PPL employs a compositional hierarchical memory, which parses poses into memory banks.
-Each memory bank explicitly contains multiple vectors encoding the variations of each sub-structure of a pose.
-This facilitates the aggregation of poses into a representative categorical pose prior.
+Recently, Hedlin et al.
+(2024) leveraged a large pre-trained text-to-image Stable Diffusion model as its prior knowledge for pose estimation.
+In contrast, our model attains comparable pose estimation accuracy while training solely on the vision modality, with substantially reduced model size.
 3 Our Proposed Pose Prior Learner (PPL)
 We introduce our proposed method, Pose Prior Learner (PPL).
-Given images featuring object instances from a specific category, such as dogs or humans, PPL can accurately estimate the poses of the objects in that category while gradually learning a general pose prior through unsupervised learning.
+Given images featuring object instances from a specific category, PPL can accurately estimate the poses of the objects in that category while gradually learning a general pose prior through unsupervised learning.
 Note that our PPL requires no extra knowledge from human annotators.
@@
 We initialize $W$ with random positive values.
-$T$ is decoded from a hierarchical memory $M$ storing compositional parts of prototypical poses, which is also randomly initialized.
-Figure 2: Overview of our proposed Pose Prior Learner (PPL).
-We first distill the keypoint prior from the hierarchical memory $M$.
-Features of the image $I$ and the embedding of the keypoint prior are concatenated to predict the affine transformation parameters.
-The keypoint prior is transformed and their pair-wise links are modulated with the connectivity prior $W$ to obtain the combined link heatmap $S$.
-The concatenation of the link heatmap $S$ and the reference image $I_{ref}$ is decoded to produce the reconstructed image $I_{recon}$.
-The $sg$ symbol represents the stopping gradient operation.
-The red arrows indicate the gradient flows during backpropagation based on image reconstruction.
-See Section 3.4 for training details.
-During training, PPL inputs an image $I$ of $\mathsf{H}\times\mathsf{W}\times 3$ where $\mathsf{H}$ and $\mathsf{W}$ are the height and the width.
-The aim of PPL is to learn to correctly predict the image-specific keypoints $T^{\prime}$ and their connectivity on $I$ from the general categorical pose prior $V=(T,W)$.
+$T$ is decoded from a hierarchical memory $M$ storing prototypical poses, which is also randomly initialized.
+The aim of PPL is to learn to correctly predict the image-specific keypoints $T^{\prime}$ and their connectivity on input image $I$ from the general categorical pose prior $V=(T,W)$.
 Ideally, if PPL makes perfect predictions of the pose on $I$, by combining it with the background information, the reconstructed image $I_{\text{recon}}$ should match $I$ exactly.
-The background information is provided by the reference image $I_{ref}$.
-For images extracted from video datasets, $I_{ref}$ can be a randomly selected frame from the same video that features the same object in a different pose.
-Alternatively, for a static image dataset, $I_{ref}$ can be a randomly masked version of the original image.
-Next, we introduce how we estimate $T^{\prime}$ on $I$ using the keypoint prior $T$.
+The background information is provided by the reference image $I_{ref}$, drawn from either image or video datasets.
+For video datasets, $I_{ref}$ corresponds to a randomly selected frame within the same video that depicts the same object in a different pose.
+For static image datasets, $I_{ref}$ is constructed by applying a random masking operation to the original image.
+Next, we introduce how to estimate $T^{\prime}$ on $I$ using the keypoint prior $T$.
 3.1 Transformation of the Keypoint Prior
@@
 The keypoint prior $T$ is converted into an embedding $h_{T}$ via a series of fully connected layers.
-Together with $h_{T}$ as inputs, PPL learns to predict the affine transformation parameters $\Theta_{i}\in[\Theta_{1},\Theta_{2},...,\Theta_{N}]$ for each keypoint $P_{i}$ in $T$ from $h_{I}$ via a two-layer fully connected network denoted as $FC(\cdot)$:
-$[\Theta_{1},\Theta_{2},...,\Theta_{i},...,\Theta_{N}]=FC(h_{I},h_{T}),$
-where $\text{where }\Theta_{i}=\begin{bmatrix}a^{(i)}&b^{(i)}&t_{x}^{(i)}\\ c^{(i)}&d^{(i)}&t_{y}^{(i)}\\ 0&0&1\end{bmatrix}.$ (1)
+Together with $h_{T}$ as inputs, PPL learns to predict the affine transformation parameters $\Theta_{i}\in[\Theta_{1},\Theta_{2},...,\Theta_{N}]$ for each keypoint $P_{i}$ in $T$ from $h_{I}$ via a two-layer fully connected network denoted as $FC(\cdot)$: $\Theta_{1},\Theta_{2},\dots,\Theta_{N}]=FC(h_{I},h_{T})$, where $\Theta_{i}=\big[\big[a^{(i)},\ b^{(i)},\ t_{x}^{(i)};\ c^{(i)},\ d^{(i)},\ t_{y}^{(i)};\ 0,\ 0,\ 1\big]\big]$.
 $t_{x}^{(i)}$ and $t_{y}^{(i)}$ are the translations and $a^{(i)},b^{(i)},c^{(i)},d^{(i)}$ are the coefficients that define rotation, scaling, and shear.
-Each point $P_{i}$ in $T$ can then be transformed by $\Theta_{i}$, resulting in the image-specific keypoints $T^{\prime}$ for the image $I$:
-$T^{\prime}=[P^{\prime}_{1},P^{\prime}_{2},...,P^{\prime}_{N}],\text{ where }[P^{\prime}_{i},1]^{\top}=\Theta_{i}[P_{i},1]^{\top}.$ (2)
+Each point $P_{i}$ in $T$ can then be transformed by $\Theta_{i}$, resulting in keypoints $T^{\prime}$ for image $I$: $T^{\prime}=[P^{\prime}_{1},P^{\prime}_{2},...,P^{\prime}_{N}],\text{ where }[P^{\prime}_{i},1]^{\top}=\Theta_{i}[P_{i},1]^{\top}$.
 3.2 Connecting Keypoints Based on the Connectivity Prior
@@
 In this section, we introduce the connectivity prior and explain how it can be used to regularize the connectivity strength between any pair of estimated keypoints in $T^{\prime}$ on $I$.
-Similar to AutoLink (He et al., 2022a), PPL connects any two keypoints $P^{\prime}_{i}$ and $P^{\prime}_{j}$ in $T^{\prime}$ to obtain differentiable link heatmap $S_{i,j}\in\mathbb{R}^{\mathsf{H}\times\mathsf{W}}$.
-Intuitively, each 2D link heatmap represents a probability density map, where the pixel values along the link between two points are high, while other areas are assigned values close to zero.
-For any point $P^{\prime}_{i}$, its strongest connectivity to any of the other points in $T^{\prime}$ is activated on the combined link heatmap $S\in\mathbb{R}^{\mathsf{H}\times\mathsf{W}}$ via a max pooling operation over all the $N\times N$ link heatmaps:
-$S=\max_{i,j}^{N\times N}(w_{i,j}S_{i,j}),$ (3)
-where $w_{i,j}$ in the connectivity prior $W$ modulates the link heatmap $S_{i,j}$ based on whether the two keypoints $P^{\prime}_{i}$ and $P^{\prime}_{j}$ are physically connected.
+Similar to AutoLink He et al.
+(2022a), PPL connects any two keypoints $P^{\prime}_{i}$ and $P^{\prime}_{j}$ in $T^{\prime}$ to obtain differentiable link heatmap $S_{i,j}\in\mathbb{R}^{\mathsf{H}\times\mathsf{W}}$.
+Each 2D link heatmap represents a probability density map, where the pixel values along the link between two points are high, while other areas are assigned values close to zero.
+For any point $P^{\prime}_{i}$, its strongest connectivity to any of the other points in $T^{\prime}$ is activated on the combined link heatmap $S\in\mathbb{R}^{\mathsf{H}\times\mathsf{W}}$ via a max pooling operation over all the $N\times N$ link heatmaps: $S=\max_{i,j}^{N\times N}(w_{i,j}S_{i,j})$, where $w_{i,j}$ in the connectivity prior $W$ modulates the link heatmap $S_{i,j}$ based on whether the two keypoints $P^{\prime}_{i}$ and $P^{\prime}_{j}$ are physically connected.
 Ideally, if PPL correctly estimates the probability of physical links for an object category, $S_{i,j}$ will receive higher connectivity values, thereby activating the locations linking these two keypoints on the combined link map $S$.
-Given the combined link map $S$ and the reference image $I_{ref}$, PPL can reconstruct the image $I$.
-$I_{ref}$ provides background information for reconstruction, while $S$ supplies foreground structural information by linking all the estimated keypoints with the connectivity prior.
-Therefore, we concatenate $S$ and $I_{ref}$ and feed them into a 2D-CNN to perform the image reconstruction $I_{recon}$, where $I_{recon}=\phi_{dec}(I_{ref},S)$.
+Given the link map $S$ and the reference image $I_{ref}$, PPL can reconstruct the image $I$.
+$I_{ref}$ provides background information for reconstruction, while $S$ supplies foreground structural information by linking the estimated keypoints with the connectivity prior.
+Therefore, we concatenate $S$ and $I_{ref}$ and feed them into a 2D-CNN to perform the image reconstruction $I_{recon}$: $I_{recon}=\phi_{dec}(I_{ref},S)$.
 3.3 Reconstruction of Keypoint Configuration with Memory
-$M$ is a memory module storing compositional representations of prototypical poses.
-$M$ is hierarchical because it consists of $m$ memory banks $\{b_{1},b_{2},...,b_{m}\}$, each containing $k$ learnable vectors of dimension $d$.
-This setup allows for efficient retrieval and management of information at multiple levels, where $M$ can represent pose configurations while memory banks capture specific pose sub-structures.
-This hierarchy enables $M$ to effectively learn robust keypoint prior representations $T$ for the following reasons: (1) By storing multiple prototypical poses, memory helps in aggregating these poses, which aids in creating a more robust and comprehensive prior that captures variations within an object category.
-(2) $M$ can assist in organizing information hierarchically, making it easier for PPL to retrieve relevant prototypical poses when making predictions in uncertain or ambugious scenarios, such as occlusion.
-and (3) By leveraging $M$, PPL can refine its predictions iteratively, using stored poses to adjust its outputs based on previously learned pose representations.
-Next, we introduce how $M$ is structured.
-Given the estimated $N$ keypoints in $T^{\prime}$, we first encode them into several tokens of dimension $d$ with several MLP-Mixer blocks (Tolstikhin et al., 2021), denoted as $MIX_{enc}$.
-The number of encoded tokens should always be equal to the number of memory banks, as we design these tokens to have different embedding space.
-Each token $g_{i}$ corresponds to the embedding of memory bank $b_{i}$ in memory $M$.
-We define $G$ as the collection of the $m$ tokens:
-$G=[g_{1},g_{2},...,g_{m}]=MIX_{enc}(T^{\prime}),$ (4)
-If all the vectors in all the memory banks of $M$ learn to capture unique parts of prototypical poses, each $g_{i}$ should be able to retrieve the most similar vector in the memory bank $b_{i}$ of $M$ and reconstruct $G$ itself.
-Here, we use $L2$ distance to measure the similarity between $g_{i}$ and $k$ vectors in $b_{i}$ of $M$.
-The vector in $b_{i}$ that is most similar to $g_{i}$ is denoted as $g^{\prime}_{i}$.
-Thus, for each $g_{i}$, we can always find the corresponding $g^{\prime}_{i}$ from $b_{i}$, resulting in a collection of $G^{\prime}=[g^{\prime}_{1},g^{\prime}_{2},...,g^{\prime}_{m}]$.
-Another series of MLP-Mixer blocks (denoted as $MIX_{dec}$) is then used to decode $G^{\prime}$ back to $N$ keypoints $T^{\prime}_{recon}$, where $T^{\prime}_{recon}=MIX_{dec}(G^{\prime})$.
-See Figure ALABEL:fig:mem_stru in Appendix A5 for the schematic.
-Unlike PCT (Geng et al., 2023) where all vectors of a memory are within the same embedding space, We organize $M$ into $m$ memory banks, each containing $k$ vectors that represent an independent embedding space.
-This structure enables us to efficiently distill each memory bank to form a general pose representation, which can be viewed as the keypoint prior $T$.
-Specifically, we define $MP(\cdot)$ as the mean pooling operation.
-PPL pools the $k$ vectors of each memory bank into one vector.
-These pooled $m$ vectors are further decoded by $MIX_{dec}$ into $N$ points of our keypoint prior $T$.
-See Figure ALABEL:fig:mem_cons in Appendix A5 for the schematic of distilling $T$: $T=MIX_{dec}([MP(b_{1}),MP(b_{2}),...,MP(b_{m})])$.
-Compared to simply using a moving average of the learned poses, the advantages of employing hierarchical memory and keypoint configuration reconstruction are as follows: (1) the hierarchical memory increases the complexity of the embedding space with compositional parts, enhances the expressive capability for poses and facilitates more accurate pose retrieval for occluded images; (2) the compositional nature of the hierarchical memory addresses data bias, enabling the distillation of a less biased prior.
+$M$ is a memory module that stores compositional representations of prototypical keypoint configurations.
+We organize $M$ hierarchically into $m$ memory banks $\{b_{1},b_{2},\dots,b_{m}\}$, each containing $k$ learnable vectors of dimension $d$.
+This hierarchical design brings three advantages: (1) Hierarchical $M$ offers richer complexity, as its capacity can scale exponentially with the number of levels.
+(2) Organizing information across banks facilitates robust retrieval of relevant prototypes in uncertain or ambiguous cases (e.g., occlusion).
+When occlusion disrupts the feed-forward visual pathway, the stored prototypical poses offer plausible hypotheses about the true pose given only partial observations.
+By iteratively refining its estimates against these prototypes, the network can “fill in” missing information and recover accurate poses, even under heavy occlusion.
+(3) Hierachical $M$ allows for efficient information retrieval at multiple levels: while $M$ encodes complete pose configurations, hirachical memory narrow down the search space by dividing $M$ into parts and individual memory banks capture finer sub-structures of poses.
+We provide a comparison with a single large memory bank in Table A3 in Appendix A8 to highlight the advantage of this hierarchical design.
+Encoding into Memory Space.
+Given the estimated keypoints $T^{\prime}$, we encode them into $m$ tokens of dimension $d$ using several MLP-Mixer blocks Tolstikhin et al.
+(2021), denoted as $MIX_{enc}$.
+The number of tokens equals the number of memory banks, with each token $g_{i}$ assigned to a distinct embedding space corresponding to memory bank $b_{i}$.
+Collectively, these tokens are $G=[g_{1},g_{2},\dots,g_{m}]=MIX_{enc}(T^{\prime})$.
+Memory Retrieval and Reconstruction.
+If the vectors in $M$ learn to represent unique parts of prototypical poses, each $g_{i}$ should retrieve its most similar vector from bank $b_{i}$.
+We measure similarity using the $L_{2}$ distance, and denote the retrieved vector as $g^{\prime}_{i}$.
+The retrieved set is $G^{\prime}=[g^{\prime}_{1},g^{\prime}_{2},\dots,g^{\prime}_{m}].$ Finally, another series of MLP-Mixer blocks ($MIX_{dec}$) decodes $G^{\prime}$ back into $N$ keypoints: $T^{\prime}_{recon}=MIX_{dec}(G^{\prime})$.
+See Figure ALABEL:fig:mem_stru in Appendix A7 for an illustration.
+Distilling the Keypoint Prior.
+Unlike PCT Geng et al.
+(2023), where all memory vectors lie in a single embedding space, we assign each bank an independent space.
+This setup allows each bank to capture distinct pose components, which can then be distilled into a general pose prior.
+Specifically, let $MP(\cdot)$ denote mean pooling.
+PPL pools the $k$ vectors within each bank into one vector, resulting in $m$ pooled vectors.
+These are decoded into $N$ keypoints by $MIX_{dec}$: $T=MIX_{dec}([MP(b_{1}),MP(b_{2}),\dots,MP(b_{m})])$.
+See Figure ALABEL:fig:mem_cons in Appendix A7 for the schematic.
+Table 1: Keypoint detection on the video datasets Human3.6m and Taichi, and the image dataset CUB-200-2011.
+For the video datasets Human3.6m and Taichi, we use video frames as $I_{ref}$ during training.
+For all the CUB evaluation subsets, we instead use randomly masked images as $I_{ref}$.
+We report mean $L_{2}$ error normalized by image resolution for Human3.6m and CUB-200-2011, and summed $L_{2}$ error for Taichi.
+We use resolution $256\times 256$ for Taichi and $128\times 128$ for CUB-200-2011.
+For Human3.6m, we report results under both resolutions.
+Following He et al.
+(2022a), we use different numbers of keypoints for the evaluation subsets of CUB-200-2011: 10 for CUB-Aligned and 4 for CUB-001, CUB-002, CUB-003, and CUB-All.
+The best results are highlighted in bold.
+A dash (–) indicates that the result of the corresponding method is not applicable in the literature.
+Method     Human3.6m     Taichi
+CUB-
+Aligned
+CUB-
+001
+CUB-
+002
+CUB-
+003
+CUB-
+All
+    Res. 128 Res. 256
+Jakab Jakab et al. (2020)     - 2.73     -     - - - - -
+Thewlis Thewlis et al. (2017)     - 7.51     -     - - - - -
+Lorenz Lorenz et al. (2019)     - 2.79     -     - - - - -
+Schmidtke Schmidtke et al. (2021)     - 3.31     -     - - - - -
+BKind Sun et al. (2022)     2.44 -     -     - - - - -
+Siarohin Siarohin et al. (2021)     - -     389.78     - - - - -
+LatentKeypointGAN He et al. (2021)     - -     437.69     5.21 22.6 29.1 21.2 14.7
+GANSeg He et al. (2022b)     - -     417.17     3.23 22.1 22.3 21.5 12.1
+Zhang Zhang et al. (2018)     - 4.91     343.67     5.36 26.9 27.6 27.1 22.4
+AutoLink He et al. (2022a)     2.76 -     316.10     3.51 20.2 19.2 18.5 11.3
+PPL (ours)     1.92 2.56     293.35     3.19 19.3 18.6 17.3 10.5
 3.4 Training and Inference
 Our PPL is trained to jointly minimize all the four losses: the image reconstruction loss $L_{ir}$, the boundary loss $L_{b}$, the link regularization loss $L_{l}$, and the keypoint configuration reconstruction loss $L_{kr}$.
-We elaborate on these four losses below.
+We describe these losses in detail below and present their ablation results in Table A3 in Appendix A8.
 Additionally, we use three training techniques for stable convergence of the network.
@@
 Image Reconstruction Loss.
-If PPL correctly estimates the pose on the original image $I$, the reconstructed image $I_{recon}$, based on the estimated pose, should be identical to $I$.
+If PPL correctly estimates the pose of image $I$, the reconstructed image $I_{recon}$, based on the estimated pose, should be identical to $I$.
 Therefore, ensuring the quality of $I_{recon}$ encourages PPL to improve its pose estimation accuracy.
-To achieve this, we apply a perceptual loss on the embeddings of $I$ and $I_{recon}$, extracted using a frozen feature extractor $\psi(\cdot)$ from the VGG19 network pre-trained on ImageNet (Russakovsky et al., 2015).
+To achieve this, we apply a perceptual loss on the embeddings of $I$ and $I_{recon}$, extracted using a frozen feature extractor $\psi(\cdot)$ from the VGG19 network pre-trained on ImageNet Russakovsky et al.
+(2015).
 The perceptual loss is defined as: $L_{ir}=\|\psi(I_{recon})-\psi(I)\|_{1}$.
 Boundary Loss.
-To ensure that the network does not transform the points in the keypoint prior outside the boundaries of the image, we limit the x and y coordinates of the transformed keypoints to be within the image:
-$L_{b}=\sum_{*\in{x,y}}\begin{cases}|P^{\prime}_{i,*}|&\text{if }|P^{\prime}_{i,*}|>1,\\ 0&\text{otherwise}.\end{cases}$ (5)
-where $P^{\prime}_{i,x}$ and $P^{\prime}_{i,y}$ are the normalized $x$ and $y$ coordinate of the keypoint $P^{\prime}_{i}$ respectively.
+To prevent the keypoint prior from being transformed outside the image boundary, we limit the transformed keypoints to be within the image: $L_{b}=\sum_{*\in\{x,y\}}(|P^{\prime}_{i,*}|\text{ if }|P^{\prime}_{i,*}|>1,\text{ else }0)$, where $P^{\prime}_{i,x}$ and $P^{\prime}_{i,y}$ are the normalized $x$ and $y$ coordinate of the keypoint $P^{\prime}_{i}$ respectively.
 Link Regularization Loss.
 A person’s arm always maintains a fixed length regardless of the poses.
-Thus, we propose the constraint that links should be assigned a high weight if they do not vary significantly in length before and after the affine transformation.
+Thus, we propose the constraint that links should be assigned a high weight if they do not vary significantly in length before and after the transformation.
+Since limb rigidity is inherently defined in 3D space, we introduce it as an auxiliary constraint in 2D that approximates this property, primarily to accelerate convergence and stabilize training.
 The loss $L_{l}$ encourages the preservation of the link lengths during pose estimation.
-It is defined as in the equation below, where $l(\cdot)$ is the $L2$ distance between two keypoints before and after the affine transformation:
-$L_{l}=\sum_{i,j}w_{i,j}\|l(P_{i},P_{j})-l(P^{\prime}_{i},P^{\prime}_{j})\|_{1}.$ (6)
+It is defined as: $L_{l}=\sum_{i,j}w_{i,j}\|l(P_{i},P_{j})-l(P^{\prime}_{i},P^{\prime}_{j})\|_{1}$, where $l(\cdot)$ is the $L2$ distance between two keypoints before and after the transformation.
 Reconstruction Loss on Keypoint Configurations.
@@
 Therefore, we introduce the keypoint configuration reconstruction loss defined as: $L_{kr}=\|T^{\prime}_{recon}-T^{\prime}\|_{2}+\|G-G^{\prime}\|_{2}$.
-Table 1: Keypoint detection on CUB-200-2011.
-We report the mean $L2$ error normalized by the image resolutions of $128\times 128$.
-We use 10 keypoints for CUB-Aligned and 4 keypoints for CUB-001, CUB-002, CUB-003, and CUB-all.
-The best is in bold.
-Method CUB- aligned CUB- 001 CUB- 002 CUB- 003 CUB- all
-(Zhang et al., 2018) 5.36 26.9 27.6 27.1 22.4
-(He et al., 2021) 5.21 22.6 29.1 21.2 14.7
-(He et al., 2022b) 3.23 22.1 22.3 21.5 12.1
-AutoLink (He et al., 2022a) 3.51 20.2 19.2 18.5 11.3
-PPL (ours) 3.19 19.3 18.6 17.3 10.5
-Table 2: Keypoint detection on Human3.6m.
-We report the mean $L2$ error normalized by the image resolutions (Res.) of both $256\times 256$ and $128\times 128$.
-* is the result trained by using masked images as reference images.
-The best is in bold.
-Method Res. Norm. $L2$ Error
-(Jakab et al., 2020) 256 2.73
-(Thewlis et al., 2017) 256 7.51
-(Lorenz et al., 2019) 256 2.79
-(Zhang et al., 2018) 256 4.91
-(Schmidtke et al., 2021) 256 3.31
-AutoLink (He et al., 2022a) 128 2.76
-PPL (ours) 128 1.92
-PPL* (ours) 128 2.23
-PPL (ours) 256 2.56
-Table 3: Keypoint detection on Taichi.
-For consistency with the baseline methods, we report the summed $L2$ error at a resolution of $256\times 256$.
-* is the result trained by using masked images as reference images.
-The best is in bold.
-Method Summed $L2$ Error
-(Siarohin et al., 2021) 389.78
-(He et al., 2021) 437.69
-(Zhang et al., 2018) 343.67
-(He et al., 2022b) 417.17
-AutoLink (He et al., 2022a) 316.10
-PPL* (ours) 298.60
-PPL (ours) 293.35
 Iterative Inference.
-We propose an iterative inference strategy (Figure 3).
-$4$ iterations are used for every experiment.
-In every iteration, we take the reconstructed image $I_{recon}$ from the last iteration (the original image $I$ for iteration 00) as the input.
+The memory in PPL stores learned prototypical poses, making it well-suited for correcting inaccurate pose estimates—especially under occlusion—through an iterative autoregressive process during inference (Figure 3).
+Iterative inference leverages this memory to progressively refine predictions alongside image reconstruction.
+In every iteration, we take the reconstructed image $I_{recon}$ from the last iteration (the original image $I$ for iteration 0) as the input.
 We infer its keypoints $T^{\prime}$ as the output keypoints of the current iteration.
@@
 We keep the original occluded image $I$ as the reference image for all the iterations.
-See Appendix A1 for more implementation details.
+$4$ iterations are used for every experiment.
+In our experiments (see Section 4.3), we demonstrate that iterative inference outperforms one-step correction.
+This advantage stems from several key factors: (1) A single correction often fails to capture complex dependencies and may introduce additional errors.
+In contrast, autoregressive inference progressively refines predictions, enhancing both accuracy and consistency.
+(2) For tasks like pose estimation, initial predictions are often ambiguous.
+Iterative updates enable the model to integrate contextual cues at each step, thereby resolving uncertainties more effectively.
+(3) Iterative inference also encourages the model to reason about structured dependencies across spatial and temporal domains, improving robustness to scene variation, object occlusion, and background noise.
 Figure 3: Overview of the iterative inference strategy in our PPL (Section 3.4).
@@
 The hierarchical memory $M$ refines the estimated pose $T^{\prime}$ and outputs $T^{\prime}_{recon}$.
-With the original image $I$ as the reference image $I_{ref}$, PPL reconstructs the image $I_{recon}$.
+The original image II is used as the reference image to reconstruct the image $I_{recon}$.
 It is then used as the input image in the next iteration.
+Training and Implementation Details.
+We use the Adam optimizer with a learning rate of $10^{-3}$, training for 50 epochs.
+The learning rate for link weights is scaled by 512 to address the small gradients of SoftPlus near zero.
+We conduct experiments using the link thickness $\sigma^{2}=5\times 10^{-4}$ across all benchmark datasets, where we adopt the same definition of $\sigma$ used in He et al.
+(2022a).
+For the hierarchical memory, we use $34$ memory banks, each of which contains $16$ vectors of dimension $512$.
+All results are averaged over 3 experiment runs.
+We train and test PPL on a single NVIDIA RTX A6000 GPU with the batch size of 64.
+Compute efficiency is reported in Table A1.
+See Appendix A1 for more details.
 4 Experiments
-For quantitative experiments, we use three image datasets: Human3.6m (Ionescu et al., 2013), Taichi (Siarohin et al., 2019b, a), and CUB-200-2011 (Wah et al., 2011).
-For fair comparisons, we set the number of keypoints $N$ to be the same as (He et al., 2022a).
-For qualitative visualizations, we use Youtube dog videos, Flowers (Nilsback & Zisserman, 2008), 11k-Hands (Afifi, 2019), and Horses (Zhu et al., 2017).
-See Appendix A2 for details.
-We utilize video frames as $I_{ref}$ for the Human3.6m, Taichi, and YouTube dog videos, while randomly masked images are used as $I_{ref}$ for other datasets.
-On Human3.6m and CUB-200-2011, we report the results in the mean $L2$ error between the predicted keypoints and the ground truth, normalized by the image size.
-For Taichi (He et al., 2022a; Siarohin et al., 2021; He et al., 2021, 2022b; Zhang et al., 2018), we use the summed $L2$ error computed at a resolution of $256\times 256$.
+For quantitative experiments, we use three public datasets: Human3.6m Ionescu et al.
+(2013), Taichi Siarohin et al.
+(2019b; a), and CUB-200-2011 Wah et al.
+(2011).
+We set the number of keypoints $N$ to be the same as He et al.
+(2022a).
+We use PPL trained on Human3.6m for all ablation studies and experiments in occluded scenes.
+For qualitative visualizations of the learned pose priors, we additionally include Youtube dog videos, Flowers Nilsback & Zisserman (2008), 11k-Hands Afifi (2019), and Horses Zhu et al.
+(2017).
+To train PPL, we utilize video frames as $I_{ref}$ on the Human3.6m, Taichi, and YouTube dog videos, while randomly masked images are used as $I_{ref}$ for other datasets.
+See Appendix A3 for dataset details.
 4.1 Unsupervised Categorical Pose Estimation
-We compare the keypoint detection results of PPL with other unsupervised pose estimation methods and present the results in Table 1 (CUB-200-2011), Table 2 (Human3.6m), and Table 3 (Taichi).
-On all datasets, PPL significantly outperforms all baselines across all image resolutions.
-Among the baselines, AutoLink (He et al., 2022a) also incorporates learnable connectivity priors.
-However, its performance is inferior due to the absence of hierarchical memory.
-Notably, even when masked images are used as $I_{ref}$, PPL consistently outperforms AutoLink.
-This demonstrates that the effectiveness of PPL is primarily attributed to its pose prior, rather than relying on cross-frame reconstruction.
-Additionally, we note that STT (Schmidtke et al., 2021), a baseline with human-defined priors, still underperforms PPL.
-Consistent with HPE (Yoo & Russakovsky, 2023), this suggests that pre-defined priors are not always optimal.
+We compare the keypoint detection results of PPL with other unsupervised pose estimation methods and present the results in Table 1.
+On all datasets, PPL outperforms all baselines across all image resolutions.
+We note that STT Schmidtke et al.
+(2021), a baseline with human-defined priors, underperforms PPL.
+Consistent with HPE Yoo & Russakovsky (2023), this suggests that pre-defined priors are not always optimal.
 PPL can learn more representative priors that outperform those manually defined.
+Among the baselines, AutoLink He et al.
+(2022a) also incorporates learnable connectivity priors; but its performance is inferior to PPL due to the absence of hierarchical memory.
+To demonstrate that PPL’s advantage comes from the learned priors rather than temporal consistency in video datasets, we treat all video frames as independent images and use randomly masked images as $I_{ref}$ to train PPL on the video datasets.
+We denote the resulting model as PPL*.
+Notably the results of PPL* on Human3.6m and Taichi in Table A4 in Appendix A8 show that even without temporal consistency in video reference frames as supervision, PPL* continues to outperform AutoLink.
+This highlights that the effectiveness of PPL stems primarily from its pose prior, rather than from cross-frame reconstruction.
+Additionally, we provide the result comparisons between our PPL and Hedlin et al.
+(2024) in Table A2 in Appendix A4.
+It is noteworthy that PPL performs competitively with Hedlin et al.
+(2024), which is based on priors from pre-trained text-image stable diffusion models, despite PPL being much smaller in size and relying solely on the visual modality rather than both text and vision.
+We also present visual comparisons of the estimated poses between our PPL and AutoLink in Figure LABEL:fig:PPL_AutoLink.
+From scenes without complex backgrounds (Row 1), both AutoLink and PPL can predict correct poses.
+However, in more challenging settings—such as those with high-contrast backgrounds (Row 2) or where arms blend into the background (Row 3)—AutoLink occasionally misplaces keypoints on the background, resulting in implausible poses.
+In contrast, PPL consistently attends to the human body, leveraging its learned pose prior to enforce realistic body constraints.
 Visualization of Pose Estimation.
-We provide the visualization results of pose estimation on Human3.6m (Figure ALABEL:human-vis), YouTube dog videos (Figure ALABEL:dog-vis), 11k-Hands (Figure ALABEL:hand-vis), CUB-200-2011 (Figure ALABEL:bird-vis), Horse (Figure ALABEL:horse-vis), and Flowers (Figure ALABEL:flower-vis) in Appendix A7.
+We provide the visualization results of pose estimation on Human3.6m (Figure ALABEL:human-vis), YouTube dog videos (Figure ALABEL:dog-vis), 11k-Hands (Figure ALABEL:hand-vis), CUB-200-2011 (Figure ALABEL:bird-vis), Horse (Figure ALABEL:horse-vis), and Flowers (Figure ALABEL:flower-vis) in Figure 7(a) in Appendix A5.
 The visualization results demonstrate that PPL can learn categorical priors and estimate poses for various object categories, without any external annotations.
@@
 In Row 2, Column 5 of Figure ALABEL:dog-vis, PPL correctly estimates the pose of a dog lowering down its head.
-Moreover, we found the quality of estimated poses of dogs is inferior to that of humans.
-This is primarily due to the significant morphological differences among various dog breeds.
-Additionally, dogs often perform actions such as turning around, which can lead to significant changes in pose that are difficult to accurately capture by our priors in 2D space.
-Additional visualization results for Flowers, Hands, and Horses in Appendix A7 further demonstrate that our PPL can be applied for various categories.
+Moreover, we found that the quality of estimated dog poses is lower than that of humans, primarily because dogs exhibit much greater variability in pose, breed, size, and fur texture.
+Additional visualization results for Flowers, Hands, and Horses in Figure 7(a) in Appendix A5 further demonstrate that our PPL can be applied for various categories.
 Notably, flowers are rigid objects with no degrees of freedom like animals.
@@
 4.2 Ablation Studies
+Ablation on Prior Variants.
+We investigate how different initializations of connectivity and keypoint priors affect pose estimation and whether further refining these priors enhances performance.
+From Table 2, we obtain several key insights: (1) Models with frozen, human-defined priors (Column 4) perform worse than PPL, indicating that PPL learns more representative priors than those predefined by humans.
+(2) Refining pre-defined keypoint and connectivity priors (Column 1) outperforms our default PPL, suggesting that PPL can enhance models with human-defined priors through refinement.
+(3) Interestingly, randomly initializing either keypoint or connectivity priors, followed by refinement during training (Columns 5-9), yields comparable performance to models with human-defined priors.
+This suggests that human-defined priors may not be necessary for effective pose estimation.
+(4) Surprisingly, freezing randomly initialized keypoint priors also results in reasonable pose estimation accuracy, though it is still lower than PPL’s default performance (Columns 7 and 9).
+(5) In contrast to (4), freezing random connectivity priors prevents the model from converging, implying that connectivity priors play a more critical role in guiding pose estimations than keypoint priors.
 Table 2: Keypoint detection results of our PPL variants on the Human3.6m dataset.
@@
 1 2 3 4 5 6 7 8 9 10 11
-Keypoint prior Initialization Pre Pre Pre Pre Pre Pre Rand Rand Rand Rand From mem
+Keypoint
+prior
+Initialization Pre Pre Pre Pre Pre Pre Rand Rand Rand Rand From mem
 Trainable
-Connectivity prior Initialization Pre Pre Pre Pre Rand Rand Pre Pre Rand Rand Rand
+Connectivity
+prior
+Initialization Pre Pre Pre Pre Rand Rand Pre Pre Rand Rand Rand
 Trainable
-Normalized $L2$ Error 2.51 2.66 2.58 2.70 2.54 2.61 2.68 2.72 2.75 2.83 2.56
-Ablation on Prior Variants.
-Here, we investigate how different initializations of connectivity and keypoint priors affect pose estimation and assess whether further refining these priors enhances performance.
-From Table 4, we obtain several key insights: (1) Models with frozen, human-defined priors (Column 4) perform worse than our PPL, indicating that PPL learns more representative priors than those predefined by humans.
-(2) Refining pre-defined keypoint and connectivity priors (Column 1) outperforms our default PPL, suggesting that PPL can enhance models with human-defined priors through refinement.
-(3) Interestingly, randomly initializing either keypoint or connectivity priors, followed by refinement during training (Columns 5-9), yields comparable performance to models with human-defined priors.
-This suggests that human-defined priors may not be necessary for effective pose estimation.
-(4) Surprisingly, freezing randomly initialized keypoint priors also results in reasonable pose estimation accuracy, though it is still lower than PPL’s default performance (Columns 7 and 9).
-(5) In contrast to (4), freezing random connectivity priors prevents the model from converging, implying that connectivity priors play a more critical role in guiding pose estimations than keypoint priors.
+Normalized
+$L2$ Error
+2.51 2.66 2.58 2.70 2.54 2.61 2.68 2.72 2.75 2.83 2.56
 Ablation on Number & Dimension of Vectors in Each Memory Bank.
@@
 Here, we analyze the impact of the number of vectors per memory bank and the dimension of each vector on PPL’s pose estimation performance.
-From Figure 5 in Appendix A3, we observed that PPL remains robust across different vector counts and dimensions, although performance slightly improves with more vectors of higher dimensions.
+From Figure A10 in Appendix A8, we observed that PPL remains robust across different vector counts and dimensions, although performance slightly improves with more vectors of higher dimensions.
 As a result, we fixed 16 vectors per memory bank, each with a dimension of 512, for all experiments.
@@
 We varied the number of keypoints in the pose priors from 4 to 32.
-The results in Figure 5 in Appendix A3 show that pose estimation accuracy improves as the number of keypoints in the prior increases.
+The results in Figure A10 in Appendix A8 show that pose estimation accuracy improves as the number of keypoints in the prior increases.
 However, using 32 keypoints offers limited improvement compared to PPL with 16 keypoints.
 4.3 Pose Estimation in Occluded Scenes
+(a) Visualization results of poses estimation on Human3.6m.
+(a) Pose estimation on occluded images in Human3.6m.
+The first column shows the original image and its estimated pose by PPL.
+Columns 2-5 show the iterative inference process where the reconstructed images by PPL (Row 1 and 3) are fed back to itself for estimating poses (Rows 2 and 4) on occluded images either using CenterMasking (Row 1 and 2) or RandomMasking (Row 3 and 4).
+(b) The pose prior evolves as a function of training epochs (from top to bottom).
+(c) Comparison between PPL and AutoLink He et al.
+(2022a) in estimated poses on example images from Human3.6m.
+The left column shows testing images, the middle column and the right column show estimated poses by AutoLink and PPL respectively.
+All results are obtained with 16 keypoints.
 To verify the robustness of PPL in occluded scenes, we divide the image into $32\times 32$ patches and apply two masking techniques: RandomMasking and CenterMasking.
@@
 Across both RandomMasking and CenterMasking, with our iterative inference strategy, PPL successfully reconstructs the occluded image parts after three iterations and meanwhile, predicts reasonable full-body poses.
+We include additional qualitative samples on occlusions in Figure A1 in Appendix A5, showing that PPL consistently outperforms AutoLink by reconstructing plausible poses even with partial visual information.
+Generally, incorrect predictions are corrected by the memory, particularly under small-area occlusions.
+However, severe occlusions can lead to plausible yet incorrect reconstructions, potentially compounding errors.
+A failure example is provided in Figure A1 in Appendix A5 where PPL fails to reconstruct the ground truth pose with large occlusion.
 5 Discussion
@@
 Extending PPL to incorporate 3D priors will be a key focus of our future research.
-6 Impact Statement
-We present the challenge of unsupervised learning of categorical priors for pose estimation and propose Pose Prior Learner (PPL) to address it.
-By leveraging learned priors, PPL enhances robustness and generalizability during inference, promoting the safe deployment of AI technologies.
-However, privacy concerns surrounding images and videos in these datasets, especially for human pose estimation, must be carefully addressed.
-Protecting personal and sensitive information remains paramount as AI advances.
 Appendix A1 Implementation Details of our Pose Prior Learner
@@
 The learning rate for link weights is scaled by 512 to address the small gradients of SoftPlus near zero.
-We conduct experiments using the link thickness $\sigma^{2}=5\times 10^{-4}$ across all benchmark datasets, where we adopt the same definition of $\sigma$ used in (He et al., 2022a).
+We conduct experiments using the link thickness $\sigma^{2}=5\times 10^{-4}$ across all benchmark datasets, where we adopt the same definition of $\sigma$ used in He et al.
+(2022a).
 For the hierarchical memory, we use $34$ memory banks, each of which contains $16$ vectors of dimension $512$, for all experiments.
-Appendix A2 Datasets
-Human3.6m (Ionescu et al., 2013) is a standard benchmark dataset for human pose estimation, consisting of 3.6 million video frames.
-These frames include both 3D and 2D keypoints and were captured in a controlled studio environment with a static background, featuring various actors.
-We adhere to the approach outlined in (Zhang et al., 2018; He et al., 2022a), focusing on six activities: direction, discussion, posing, waiting, greeting, and walking.
-For training, we use subjects 1, 5, 6, 7, 8, and 9, and for testing, we use subject 11.
-Taichi (Siarohin et al., 2019b, a) consists of 3,049 training videos and 285 test videos featuring individuals performing Tai-Chi, with diverse foreground and background appearances.
-Following the approach in (Siarohin et al., 2021), we use 5,000 frames for training and 300 frames for testing.
-YouTube dog videos are videos with green backgrounds collected from YouTube to further qualitatively demonstrate the performance of PPL.
-Existing dog datasets are not suitable, as they often contain multiple, partially occluded dog instances.
-Therefore, we curated a custom dataset using 20 YouTube videos, extracting 2,000 frames for training.
-This dataset allows us to demonstrate PPL’s ability to learn pose priors for non-human categories without using ground truth poses.
-All images are trained and tested at a resolution of $256\times 256$.
-We use $10$ keypoints for this category and provide the visualization of the estimated poses on the test videos.
-The dataset will be publicly released together with other data, models, and source code.
-CUB-200-2011 (Wah et al., 2011) contains 11788 images of birds.
-We crop and align the images according to (He et al., 2022a).
-We use the train/val/test split of (Choudhury et al., 2021).
-All images are trained and tested at a resolution of $128\times 128$.
-11k-Hands (Afifi, 2019), Horse (Zhu et al., 2017), and Flowers (Nilsback & Zisserman, 2008) are used for qualitative visualization.
-Images with multiple horses are removed and all horses are aligned to face left.
-All images are trained and tested at a resolution of $128\times 128$.
-Appendix A3 Ablation on Memory Bank Vectors and Keypoint Numbers
-Figure A0: Ablation of our PPL method on memory bank vectors and number of keypoints in Human3.6m.
-The upper horizontal axis is the number of keypoints (ranging from 4 to 32) and the lower horizontal axis is the dimension of memory bank vectors (ranging from 64 to 512).
-The dashed purple line is for ablations on number of keypoints and the solid lines are for ablations on memory bank vectors.
-Appendix A4 PPL Results of Keypoint Detection on Occluded Images
-Figure A1: PPL results of keypoint detection as a function of number of inference iterations on images with RandomMasking from Human3.6m.
-The “mask ratio” in the legend specifies the masked proportion on the $32\times 32=1024$ image patches.
-Figure A2: PPL results of keypoint detection as a function of number of inference iterations on images with CenterMasking from Human3.6m.
-The “mask size” in the legend refers to the width and height of the masked region, on $1024$ image patches.
-Appendix A5 Retrieval and Distillation of the proposed Hierarchical Memory in PPL
-Appendix A6 Training Techniques
-To ensure convergence and stability during the training of PPL, we introduce three gradient dettachment techniques: (1) To address the broken gradient issue during the quantization step from $G$ to $G^{\prime}$, we adopt the approach from VQ-VAE (Van Den Oord et al., 2017).
+We utilize video frames as $I_{ref}$ for the Human3.6m, Taichi, and YouTube dog videos, while randomly masked images are used as $I_{ref}$ for other datasets.
+On Human3.6m and CUB-200-2011, we report the results in the mean $L2$ error between the predicted keypoints and the ground truth, normalized by the image size.
+For Taichi He et al.
+(2022a); Siarohin et al.
+(2021); He et al.
+(2021; 2022b); Zhang et al.
+(2018), we use the summed $L2$ error computed at a resolution of $256\times 256$.
+We excluded the DeepFashion Liu et al.
+(2016), CelebA Liu et al.
+(2015), and Zebra Zhu et al.
+(2017) datasets used by AutoLink because they either lack structural or background complexity, offer limited pose variation, or are redundant with existing categories.
+All results are averaged over 3 experiment runs.
+We train and test PPL on a single NVIDIA RTX A6000 GPU with the batch size of 64.
+The training and inference efficiency are presented in Table A1.
+Table A1: Training and inference efficiency of PPL on a single NVIDIA RTX A6000 GPU with batch size of 64.
+Resolution Training Time Inference Speed Inference Speed GPU Memory
+(per batch) (0 iteration, img/s) (3 iterations, img/s) (usage)
+$\times128$ $\sim$1.3 sec $\sim$145 $\sim$64 $\sim$9.2 GB
+$\times256$ $\sim$4.0 sec $\sim$67 $\sim$24 $\sim$37.0 GB
+Appendix A2 Training Techniques of our Pose Prior Learner
+To ensure convergence and stability during the training of PPL, we introduce three gradient dettachment techniques: (1) To address the broken gradient issue during the quantization step from $G$ to $G^{\prime}$, we adopt the approach from VQ-VAE Van Den Oord et al.
+(2017).
 Specifically, our PPL copies the gradients of $G^{\prime}$ to $G$ for backward propagation, allowing the gradients to flow through the quantization step.
@@
 Second, we detach the gradients from $T^{\prime}$ to train the memory encoder and decoder, $MIX_{\text{enc}}$ and $MIX_{\text{dec}}$, as indicated by the green arrows in Figure ALABEL:fig:mem_stru in Appendix A7.
-Appendix A7 Additional Visualizations
+Appendix A3 Datasets
+Human3.6m Ionescu et al.
+(2013) is a standard benchmark dataset for human pose estimation, consisting of 3.6 million video frames.
+These frames include both 3D and 2D keypoints and were captured in a controlled studio environment with a static background, featuring various actors.
+We adhere to the approach outlined in Zhang et al.
+(2018); He et al.
+(2022a), focusing on six activities: direction, discussion, posing, waiting, greeting, and walking.
+We use subjects 1, 5, 6, 7, 8, and 9 for training; and we use subject 11 for testing.
+Taichi Siarohin et al.
+(2019b; a) consists of 3,049 training videos and 285 test videos performing Tai-Chi, with diverse foreground and background appearances.
+Following the approach in Siarohin et al.
+(2021), we use 5,000 frames for training a linear regression and 300 frames for testing.
+YouTube dog videos are videos with green backgrounds collected from YouTube to further qualitatively demonstrate the performance of PPL.
+Existing dog datasets are not suitable, as they often contain multiple, partially occluded dog instances.
+Therefore, we curated a custom dataset using 20 YouTube videos with a sampling frequency of 30 frames per second.
+We manually curate 2,000 training frames by selecting those frames with clear dog poses and removing redundant or similar ones.
+All videos in our custom dataset are publicly available, and are used strictly for research purposes in compliance with applicable legal and ethical guidelines.
+This dataset allows us to demonstrate PPL’s ability to learn pose priors for non-human categories without using ground truth poses.
+All images are trained and tested at a resolution of $256\times 256$.
+We use $10$ keypoints for this category and provide the visualization of the estimated poses on the test videos.
+The dataset will be publicly released together with other data, models, and source code.
+CUB-200-2011 Wah et al.
+(2011) contains 11788 images of birds across 200 bird species with detailed annotations such as bounding boxes, part locations, and attributes.
+We crop and align the images according to He et al.
+(2022a).
+We use the train/val/test split of Choudhury et al.
+(2021).
+The dataset supports both aligned and non-aligned evaluation protocols, where aligned evaluation uses cropped bird regions while non-aligned evaluation relies on the original full images, making the task more challenging.
+The aligned subset CUB-Aligned and non-aligned subsets CUB-001/002/003/all are used for evaluation.
+All images are trained and tested at a resolution of $128\times 128$.
+11k-Hands Afifi (2019), Horse Zhu et al.
+(2017), and Flowers Nilsback & Zisserman (2008) are used for qualitative visualization.
+Images with multiple horses are removed and all horses are aligned to face left.
+All images are trained and tested at a resolution of $128\times 128$.
+Appendix A4 Additional Comparison with Methods Using Multi-modal Knowledge
+To provide broader context, we further compare our method with Hedlin et al.
+(2024).
+Hedlin et al.
+(2024) requires a Conditional Stable Diffusion Model pre-trained with large-scale multimodal data (image and text) to start with, thereby leveraging knowledge from an additional modality (text).
+The comparison results are presented in Table A2.
+Compared with Hedlin et al.
+(2024), our PPL method outperforms this baseline on Human3.6m and achieves comparable results on Tai-Chi for human pose estimation.
+For bird pose estimation, PPL achieves stronger performance on the CUB aligned evaluation set but lags behind on the non-aligned set.
+Overall, PPL employs a substantially smaller model, yet remains competitive with multimodal methods trained with extensive language supervision.
+Table A2: Comparison with Hedlin et al.
+(2024) on keypoint detection.
+For Human3.6m and Taichi, we report the mean L2 error normalized by the image resolution of 128×128 and the summed L2 error at a resolution of 256×256, respectively.
+For CUB-200-2011, we report the mean L2 error normalized by the resolution of 128×128.
+We use different numbers of keypoints for the subsets of CUB-200-2011: 10 for CUB-Aligned and 4 for CUB-001, CUB-002, CUB-003, and CUB-All.
+The best results are in bold.
+Human3.6m Taichi CUB-Aligned CUB-001 CUB-002 CUB-003 CUB-all
+Hedlin et al. (2024) 4.45 234.89 5.06 10.5 11.1 10.3 5.4
+PPL (ours) 1.92 293.35 3.19 19.3 18.6 17.3 10.5
+Appendix A5 Additional Visualization Results of Poses under Occlusion and Priors from Other Categories
+Figure A0: Comparison between PPL and AutoLink pose estimation results on occluded images on Human3.6m.
+The left column shows testing images, the mid column shows results of AutoLink, and the right column shows results from PPL.
+All results are obtained with 16 keypoints.
+Figure A1: One example where PPL fails to estimate the right pose when there is a large occlusion.
+From left to right, we show the original unoccluded image, the largely occluded image, the result without iterative inference (iteration 0), and the final result with iterative inference (iteration 3).
+(a) Additional visualization on (a) Human3.6m, (b) Youtube dog videos, (c) 11k-Hands, (d) CUB-200-211, (e) Horse, and (f) Flowers.
+Columns 1-4 of every subfigure show the original images and corresponding pose estimation results by PPL.
+The column 5 of every subfigure shows the learned prior for the category.
+Appendix A6 PPL Results of Keypoint Detection on Occluded Images
+Figure A7: PPL results of keypoint detection as a function of number of inference iterations on images with RandomMasking from Human3.6m.
+The “mask ratio" in the legend specifies the masked proportion on the $32\times 32=1024$ image patches.
+Figure A8: PPL results of keypoint detection as a function of number of inference iterations on images with CenterMasking from Human3.6m.
+The “mask size" in the legend refers to the width and height of the masked region, on $1024$ image patches.
+Appendix A7 Retrieval and Distillation of the proposed Hierarchical Memory in PPL
+(a) Retrieval and distillation of the proposed hierarchical memory in our PPL.
+(a) The hierarchical memory $M$ is trained to reconstruct the keypoints $^{\prime}_{recon}.
+TT^{\prime}$ is encoded into $m$ tokens by the MLP-Mixer blocks $MIX_{enc}$.
+Each token $g_{i}$ retrieves its closest vector $g^{\prime}_{i}$ in memory bank $b_{i}$.
+The resulting $m$ vectors are decoded by the MLP-Mixer $MIX_{dec}$ into the reconstructed keypoints $T^{\prime}_{recon}$.
+The green arrows indicate the gradient flows during backpropagation based on the reconstruction of keypoint configurations.
+See Section 3.4 for training details.
+(b) The hierarchical memory $M$ is distilled into the keypoint prior $T$.
+Vectors in every memory bank $b_{i}$ are mean-pooled into one vector, and the resulting $m$ vectors are decoded by $MIX_{dec}$ into the keypoint prior $T$.
+See Section 3.3 for details.
+Appendix A8 Additional ablation results
+Number of memory bank vectors and keypoints in Human3.6m.
+We vary the dimensionality of the vectors in each memory bank to examine its impact on pose estimation accuracy.
+In addition, we analyze how the number of keypoint priors influences performance.
+Results are shown in Figure A10.
+Figure A10: Ablation of our PPL method on memory bank vectors and number of keypoints in Human3.6m.
+The upper horizontal axis is the number of keypoints (ranging from 4 to 32) and the lower horizontal axis is the dimension of memory bank vectors (ranging from 64 to 512).
+The dashed purple line is for ablations on number of keypoints and the solid lines are for ablations on memory bank vectors.
+Ablations on loss components.
+Our method employs four loss components: Image Reconstruction Loss, Boundary Loss, Link Regularization Loss, and Keypoint Configuration Reconstruction Loss.
+The Image Reconstruction Loss provides the primary training signal, ensuring that the learned keypoints are semantically meaningful and spatially aligned with the object.
+The Keypoint Configuration Reconstruction Loss is critical for the successful training of the hierarchical memory.
+The Boundary Loss and the Link Regularization Loss are auxiliary constraints that mainly stabilize and accelerate convergence.
+Without the Boundary Loss, the network becomes unstable during early training.
+Replacing it with a sigmoid function causes a clear performance drop, increasing the normalized L2 error from 2.56 to 2.70 on Human3.6M at resolution 256×256.
+Similarly, removing the Link Regularization Loss leads to a measurable decline, with the error rising from 2.56 to 2.64.
+The comparison results are shown in Table A3.
+Table A3: Ablation studies on Human3.6M at resolution $256\times 256$. Numbers indicate normalized L2 error (lower is better). Best is in bold.
+Ablation Results Normalized L2 Error
+PPL-full (ours) 2.56
+w/o Boundary Loss training unstable
+Boundary Loss $\to$ Sigmoid 2.70
+w/o Link Regularization Loss 2.64
+PPL-1MemBank 2.62
+Removal of hierarchy in the memory.
+We introduce an ablated variant of PPL using only one memory bank, referred to as PPL-1MemBank.
+Below, we introduce its implementation.
+If each pose’s keypoints were encoded into 34 vectors sharing the same embedding space, mean pooling these embeddings would not effectively distill the keypoint prior.
+Instead, PPL-1MemBank encodes all keypoints of a pose into a single vector, ensuring that all vectors share a common embedding space.
+To maintain comparable capacity, PPL-1MemBank uses a single memory bank with 512 vectors, equivalent to PPL’s 34 memory banks with 16 vectors each.
+The keypoint prior is then obtained by mean pooling across this single memory bank.
+We evaluated PPL-1MemBank against PPL, with results summarized in the last row of Table A3.
+On the Human3.6m dataset, PPL-1MemBank achieves a normalized L2 error of 2.72, compared to 2.56 with PPL’s multi-bank approach.
+This indicates that multiple smaller memory banks improve pose estimation performance.
+PPL trained with randomly masked frames as reference images on video datasets.
+We also train PPL on Human3.6m and Taichi by treating video frames as independent images and using randomly masked images as $I_{ref}$.
+We denote this setting as PPL*.
+The results are reported in Table A4.
+Notably, even with masked images as $I_{ref}$, PPL* consistently outperforms AutoLink.
+This highlights that the effectiveness of PPL stems primarily from its pose prior, rather than from cross-frame reconstruction.
+Table A4: Results of Training PPL* using the randomly selected and masked video frames as $I_{ref}$ on Human3.6M and Taichi video datasets.
+We use resolution 256×256 for Taichi and 128×128 for Human3.6m.
+We use the evaluation metrics as reported in Table 1.
+Best is in bold.
+Methods Human3.6m Taichi
+AutoLink 2.76 316.10
+PPL (ours) 1.92 293.35
+PPL* (ours) 2.23 298.60